### v2

# Shamier Settle, Matthew Shapiro, Emerita Torres, and Anand Vidyashankar for discussion and feedback on the work described

Abstract
We estimate the number of street vendors in New York City. First, we summarize the process by which vendors receive licenses and permits to operate legally in New York City. Second, we describe a survey that was administered by the Street Vendor Project while distributing Coronavirus relief aid to vendors operating in New York City both with and without a license or permit. Third, we calculate the total number of vendors using ratio estimation, finding approximately 23,000 street vendors operate in New York City—20,500 mobile food vendors and 2,300 general merchandise vendors—with one third located in just six ZIP Codes—11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan. Fourth, we provide a theoretical justification of our estimates based on the theory of point processes. Finally, we evaluate the accuracy of the ratio estimator when the distribution of vendors is explained by a Poisson or Yule process, and we discuss several policy implications. In particular, our estimates suggest the American Community Survey misses the majority of New York City street vendors.
1. Introduction
Street vendors are New York City’s smallest businesses, selling food and merchandise from carts, stalls, and trucks throughout the five boroughs. They are an iconic part of the urban landscape and a thriving sector of the local economy, contributing millions of dollars in government revenue through taxes, fines, and fees. [^1] Fees from mobile food vending licenses and permits provide nearly a million dollars in revenue to New York City each year (Mosher and Turnquist 2024). Fines paid by vendors provide approximately \$200,000 each year according to data from the New York City Office of Administrative Trials and Hearings. Perhaps most importantly, street vending historically benefits underserved communities, both because vendors operate in neighborhoods with limited access to traditional stores and because vending is one of a handful of occupations in which New Yorkers of all backgrounds, immigrants in particular, are able to achieve economic mobility and a chance at the American dream (Burrows and Wallace 1998, chap. 42).
Yet despite their social and economic importance, little is known about the size and location of New York City’s street vending population. This is because while local law requires street vendors to obtain licenses and permits to operate legally, the number of licenses and permits are limited, resulting in a largely unknown population of vendors that operate without a license or permit. These vendors are not easily identified from administrative datasets, such as tax records or fines, and they can be difficult to locate for government surveys, such as the American Community Survey. Nevertheless, understanding the size and location of New York’s street vending population is crucial for informing policy and advocacy.
In this paper, we propose an estimate for the number of street vendors in New York City, including those that operate without a license or permit. We present our work in four sections. In Section 2, we review the process by which vendors receive permits and licenses, and we describe a survey administered by the Street Vendor Project at the Urban Justice Center while distributing Coronavirus relief aid to vendors operating both with and without a license or permit.
In Section 3, we use the fact that the number of licenses and permits are limited by law to construct a ratio estimator for the number of street vendors. We find approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,300 general merchandise vendors. One third are located in just six ZIP Codes: 11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan. We then provide a theoretical justification of our estimates, assuming the spatial distribution of survey respondents is representative and well-approximated by a family of inhomogeneous Poisson processes.
In Section 4, we examine the validity of these assumptions. We evaluate whether the spatial distribution of respondents is representative by comparing it to the American Community Survey and administrative data, and we evaluate the Poisson process assumption by considering a more general model in which the arrival of street vendors follows a pure birth process, of which the Poisson and Yule processes are special cases.
In Section 5, we discuss some implications of our estimates. In particular, we find that though the spatial distribution of respondents matches the results of the American Community Survey, the two disagree on the size of New York City’s street vending population. Our estimates suggest the American Community Survey misses the majority of New York City street vendors.
2. We consider two types of vendors: food vendors and merchandise vendors
A street vendor is any individual who sells goods from a mobile vending unit instead of a store. We distinguish between the mobile vending unit from which the goods are sold (i.e., the vending establishment) and the individuals who own and/or operate that unit (i.e., the vendors). Note that an establishment refers to a single vending unit. There may be multiple vendors associated with any establishment, and multiple establishments may be owned by a single individual or firm. [^2] For example, suppose Sally owns the business Sally’s Salads, which consists of two food trucks, each staffed by three employees. Then Sally’s Salads counts as one business, two establishments, and seven vendors.
We limit our analysis to two types of street vendors: mobile food vendors—vendors who sell food items such as sandwiches, drinks, and cut fruit—and general merchandise vendors—vendors who sell merchandise items such as electronics, clothing, and accessories. We refer to these vendors as food vendors and merchandise vendors, respectively. In Section 2.1, we review the process by which vendors receive permits and licenses as required by law. In Section 2.2, we describe a survey administered by the Street Vendor Project while distributing Coronavirus relief aid to vendors operating in New York City both with and without a license or permit.
2.1 The number of licenses and permits are limited by local law
A long list of rules determines the individuals, locations, and time periods during which vendors can legally sell their goods in New York City. Most relevant is the fact that a merchandise vendor requires a license to operate legally, and a food vendor requires both a license and a permit. The number of food permits and merchandise licenses are limited by law: 5,100 permits and 853 general merchandise licenses are available. Mobile food vending licenses are not limited. Note that a mobile food vending permit is issued to an individual or business to allow for the sale of food from a specific mobile food vending unit, such as a cart or truck. A mobile food vending license authorizes an individual to prepare or serve food from a permitted mobile food vending unit. [^3] Note that our estimate reflects the number of street vendors as of 2021, prior to the implementation of Local Law 18 of 2021, which was enacted concurrently. Local Law 18 made several changes to the license and permit process. These changes do not impact our estimates and therefore are not discussed in this paper.
There are several types of mobile food vendor permits with varying restrictions. Of the 5,100 permits available, 200 are borough permits that limit vendors to one of the five boroughs; 100 are reserved for veterans or vendors with a disability; 1,000 are seasonal and valid only from April to October; 1,000 are green cart permits that limit vendors to selling fruit, vegetables, plain nuts, and water; and 2,800 are unrestricted. Multiple individuals with food vendor licenses can legally operate from one permitted food vending unit. Merchandise licenses are renewed annually, while food licenses and permits are renewed biennially.
There are three relevant exceptions to these rules. The first exception is that merchandise vendors who are veterans are not subject to the limit of 853 licenses. According to data obtained by Mosher and Turnquist (2024), there are approximately 1,000 licensed merchandise vendors who are veterans in New York City. We assume the number of veterans selling merchandise without a license is negligible.
The second exception is First Amendment vendors. A First Amendment vendor sells expressive merchandise such as newspapers, books, and art. Expressive merchandise is considered free speech and protected under the First Amendment of the U.S. Constitution so that the number of First Amendment vendors cannot be restricted by law.
The final exception are concessionaires that operate on New York City parkland through an NYC Parks permit or license. According to the New York City Department of Parks and Recreation, there are approximately 400 concessions within New York City parks. Many offer food services, ranging from food carts to restaurants. We do not consider First Amendment vendors or park concessionaires in this paper. Our estimates are limited to food and merchandise vendors operating outside of parks.
2.2 The Street Vendor Project surveyed two thousand street vendors in 2021
The Street Vendor Project (SVP), part of the Urban Justice Center, is a non-profit organization that advocates on behalf of New York City street vendors. SVP administered a survey to approximately 2,000 street vendors while distributing Coronavirus relief aid in 2021. The aid was a one-time payment of \$1,000, available to any individual that owned, operated, or was otherwise employed by a street vending business in New York City between 2020 and 2021. First Amendment vendors were also eligible. There were no limits based on the size of the business, the number of sales, or whether licensed and/or permitted. All individuals eligible for aid were invited to complete a survey.
The population of street vendors estimated in this paper is the population eligible for aid (as determined by SVP) that self-identifies as either a food or merchandise vendor. SVP found aid-eligible individuals through membership lists, referrals, and canvassing operations in which SVP affiliates visited neighborhoods. Survey operations continued until \$2,415,000 in funds were distributed, yielding 2,060 responses.
The survey included 100 questions and was conducted in eight different languages: Arabic, Bangla, Cantonese, English, French, Mandarin, Spanish, Tibetan, and Wolof. [^4] Vendors who spoke other languages, such as Dari, Farsi, Russian, and Turkish, typically spoke English with a high level of fluency, and those surveys were conducted in English. Similarly, vendors who spoke indigenous languages from Latin America, such as Nahuatl or Quechua, often spoke Spanish with a high level of fluency. The survey items solicited a variety of information from vendors, such as logistical information (e.g., vending location, residential location, and frequency of operation), economic information (e.g., items sold, income, and expenses), and demographic information (e.g., age, race, ethnicity, and immigration status). Most relevant is the fact that respondents classified themselves by the goods they sold (e.g., food vendors, merchandise vendors, First Amendment vendors, etc.), and respondents indicated whether they had the relevant licenses and permits to vend. Of the 2,060 responses, 1,400 identified as food vendors and 559 as merchandise vendors. The remaining 101 respondents were predominantly First Amendment Vendors, which we exclude from our analysis.
Of the 1,400 food vendors, 349 (25%) indicated they had a permit to vend. Of the 559 merchandise vendors, 505 were not veterans, of which 308 (61%) indicated they had a license to vend. The number of respondents is listed by neighborhood in Section 7.1, Table 1. (Neighborhoods with few respondents are grouped together.) A map of the number of respondents by ZIP Code Tabulation Area (ZIP Code) is shown in the top left panel of Section 7.2, Figure 1. Approximately five percent of vendors did not identify their vending location. These vendors are excluded from Table 1 and Figure 1, although doing so does not meaningfully change the estimated number of vendors that operate in New York City. Veteran merchandise vendors are also excluded from Table 1 and Figure 1, although these vendors are reflected in our city-wide estimate of 23,000 vendors.
3. We calculate the number of street vendors using ratio estimation
We estimate approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,300 general merchandise vendors. We arrive at the estimate by using ratio estimation, which leverages the fact that the number of licenses and permits are limited by local law.
In Section 3.1, we provide a simple explanation of ratio estimation, which is intended to be accessible to a general readership. In Section 3.2, we provide a theoretical justification of our estimates based on the theory of point processes, which, while more technical, is intended to provide a somewhat general argument for ratio estimation with spatial data when a design-based approach cannot be justified. Inference is conducted using increasing domain asymptotics, and the details are outlined in Section 7.3.
3.1 We estimate 23,000 street vendors operate in New York City
Ratio estimation is a common approach for estimating the size of a population. See Cochran (1978) and Hald (1998, chap. 16) for a historical discussion and Lohr (2021, chap. 4) for an introduction. The popularity of ratio estimation is due in large part to its simplicity: Among the $1,400$ food vendors surveyed, roughly a quarter had a permit. Suppose this ratio is representative in that one quarter of all food vendors have a permit. It follows that the total number of food vendors is $20,500\approx 5,100/(1/4)$ because one quarter of $20,500$ is approximately $5,100$, the number of permits available.
By similar reasoning—i.e., exploiting the fact that the number of general merchandise licenses is limited for non-veterans—the number of non-veteran merchandise vendors is estimated to be approximately $1,300$. When combined with the estimated $1,000$ veteran merchandise vendors (see Section 2.1) and $20,500$ food vendors, the total number of street vendors is approximately $23,000$.
The main assumption of ratio estimation is that the survey data are representative in the sense that, were survey operations to continue, the percentage of respondents with a permit (in the case of food vendors) or a license (in the case of merchandise vendors) would remain the same. If we further assume that the spatial distribution of respondents is representative, we can also estimate the total number of vendors by neighborhood. [^5] We use the same ratio for each neighborhood. Ratio estimation by neighborhood is not possible since only the locations of the vendors who participated in the survey are known. These estimates are listed in Section 7.1, Table 1. We find that approximately one quarter of street vendors are located in the Chelsea and Clinton neighborhoods of Manhattan. Another quarter is located in West Queens and Lower Manhattan.
We check our estimates by comparing them to a second, independent assessment conducted by the Street Vendor Project in the Bronx on May 13 and 15, 2022, in which the number of visible street vendors was documented through on-the-ground observations. The independent assessment identified 188 total street vendors in the Fordham area (Fordham Road BID and Street Vendor Project 2024), whereas Table 1 indicates 206 vendors in Bronx Park and Fordham—an estimate approximately ten percent higher. To demonstrate that this difference can be explained by sampling variation, we derive the standard error of the ratio estimate in Section 3.2.
3.2 We model the location of street vendors as a Poisson process
We provide a point process justification of the ratio estimation procedure described in Section 3.1. The purpose is to examine the underlying assumptions and derive the standard error of the estimate.
Let $\{\Pi_{i}\}$ denote a family of inhomogeneous spatial Poisson processes, referencing the location $x\in A\subset\mathbb{R}^{2}$ of each street vendor at the time the survey was conducted. The index $i$ denotes the status of the vendor. To fix ideas, let $i=0$ indicate mobile food vendors with a permit, and let $i=1$ indicate mobile food vendors without a permit.
The Poisson process assumption may be justified by the law of rare events. New York City can be partitioned into a large number of theoretically vendable locations. Whether a vendor is located within a partition has a vanishingly small probability, such that the number of vendors in any area is well approximated by a Poisson distribution. This approximation is accurate even if these probabilities are weakly dependent. For example, see Freedman (1974), Chen (1975), and Serfling (1975). Note that other justifications are possible. For example, see Section 4.2 for a justification of the Poisson process based on the theory of birth processes.
[…]
In Section 3.1, we described the main assumption of ratio estimation: that the status and spatial distribution of the respondents is representative. We now make this statement precise. We assume that $p_{i}(x)$ can be decomposed into a constant that does not depend on $i$ or $x$ plus a spatially varying error term, $p_{i}(x)=p\,+\,\epsilon_{i}(x)$, and the error term is orthogonal to the corresponding mean measure. i.e.,
[…]
It follows that
[…]
so that by conditioning on $N_{1}(A)$, we arrive at the following probability model for the number of respondents with and without permits
$n_{0}(A)\sim$ $\ \text{Poisson}\big{(}p\,\theta\big{)}$
$n_{1}(A)\ |\ N_{1}(A)\sim$ $\ \text{Binomial}\big{(}p,\,N_{1}(A)\big{)}$
The maximum likelihood estimates for $p$ and $\theta$ are
$\hat{p}=\dfrac{n_{1}(A)}{N_{1}(A)}\qquad\text{and}\qquad\hat{\theta}=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}$
The estimator $\hat{\theta}$ is called the ratio estimator because it depends on the ratio of random variables $n_{0}(A)\,/\,n_{1}(A)$. (The total number of permitted vendors, $N_{1}(A)$, is considered fixed.) Using increasing domain asymptotics, the estimator is asymptotically normal with mean $\theta$ and standard error
$\text{SE}(\hat{\theta})=\theta\sqrt{\dfrac{1}{p\,\theta}+\dfrac{1-p}{p\,N_{1}(A)}}$
Substituting $\hat{p}$ and $\hat{\theta}$ for $p$ and $\theta$ yields the following plug-in estimator for the standard error,
$\hat{\text{SE}}(\hat{\theta})=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(A)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
See Section 7.3 for details. The margin of error reported in Section 7.1, Table 1 is two standard errors, half of the width of a 95% confidence interval. Note that the margin of error for the 206 vendors estimated in Bronx Park and Fordham is 44, and thus this estimate is consistent with the 188 street vendors counted independently by Fordham Road BID and Street Vendor Project (2024) as described in Section 3.1.
4. We examine the validity of our assumptions using additional information
We make two assumptions in Section 3. The main assumption is that the status and spatial distribution of the respondents is representative. This assumption is necessary to ensure the ratio estimator is consistent. A secondary assumption is that the spatial distribution of vendors is well described by a family of inhomogeneous spatial Poisson processes. This assumption is necessary to derive the standard errors.
In this section, we examine each assumption. Since neither can be evaluated from the survey data alone, we consider additional information. In Section 4.1, we compare the spatial distribution of the respondents to data from the American Community Survey Public Use Micro Sample 2017-2021 and the New York City Office of Administrative Trials and Hearings. In Section 4.2, we model the arrival of street vendors as a pure birth process, and we derive a more general formula for the standard errors in which the amount of extra-Poisson variation can be determined by the number of vendors that cluster within markets. [^6] The results of this section can also be derived from a spatial Negative Binomial Lévy Process in a manner that mirrors Section 3.2. The advantage of the birth process approach taken in Section 4.2 is that it provides a plausible explanation for why the Poisson assumption made in Section 3.2 may not hold in practice.
4.1 The Street Vendor Project survey is spatially representative
The main assumption in Section 3 is that the status and spatial distribution of the respondents is representative (i.e., the response probability is spatially uncorrelated with the mean measure). If this assumption holds, then the ratio estimator is consistent for a wider class of point processes in which Campbell’s Theorem applies, not just the Poisson process. See Daley and Vere-Jones (2007) for a general statement of Campbell’s Theorem.
If this assumption is violated, then $\hat{\theta}$ may be inconsistent. For example, if respondents are more likely to come from locations in which a higher proportion of vendors have permits, then the ratio $n_{0}(A)\,/\,n_{1}(A)$ would not be representative, and $\hat{\theta}$ would likely underestimate $\theta$ even in large samples.
We think this assumption is reasonable because the Street Vendor Project canvassed New York City to distribute a large amount of relief aid. We argue that trust in SVP and its mission, along with the aid, removed many barriers that typically prevent survey operations from reaching hard-to-reach populations. Instead, nonresponse reflects chance variation in the personal circumstances of a vendor that are largely unrelated to location.
Nevertheless, it is possible that canvassing operations systematically missed locations, and these missed locations may be revealed by examining other data sources. We thus compare the spatial distribution of vending locations reported by respondents to the spatial distribution of vendors who allegedly violated street vending laws during the year 2021. We consider the location of an individual’s first violation in 2021, reported in administrative data publicly available from the New York City Office of Administrative Trials and Hearings (OATH). We also compare the residencies reported by the respondents to the residencies of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers (Standard Occupational Classification 41-9091) estimated in the five year American Community Survey Public Use Micro Sample, 2017-2021 (ACS).
These comparisons are visualized in Section 7.2, Figure 1. We find the spatial distribution of vending locations largely agree (top panels) as do the spatial distribution of the residencies (bottom panels), although there are several small discrepancies. For example, the OATH data suggest a larger percentage of vendors work in South Brooklyn (Coney Island and Brighton Beach) than captured by the Street Vendor Project survey, and the ACS data suggest a larger percentage of street vendors live in South Queens (Rockaway). However, both the OATH and ACS data suggest these locations reflect only a small percentage of New York City vendors, and thus any resulting undercount is likely to have little impact on our overall estimate.
We also note that even if this discrepancy suggests a location was missed by the Street Vendor Project, our main assumption is not necessarily violated. That is because this undercount may be compensated by an overcount in other areas (e.g., the Upper East Side of Manhattan or East Brooklyn) such that our overall estimate is still accurate.
4.2 We model the arrival of street vendors as a pure birth process
The second assumption in Section 3.2 is that the spatial distribution of vendors is well approximated by a family of inhomogeneous Poisson processes. We consider this assumption secondary because the consistency of $\hat{\theta}$ holds under the wide class of point processes described in Section 4.1. That said, the standard errors derived in Section 3.2 do not necessarily hold when the Poisson assumption is violated. [^7] The ratio estimator may no longer be asymptotically efficient as well. i.e., In Section 7.3, we show that under the Poisson process assumption, the ratio estimator maximizes the likelihood, which implies the ratio estimator is asymptotically efficient. However, the ratio estimator may not maximize the likelihood when the Poisson process assumption is violated.
We think the Poisson assumption is reasonable because a wide class of data generating processes are well-approximated by the Poisson process under the law of rare events as discussed in Section 3.2. Nevertheless, to evaluate how the standard errors might change when this assumption is violated, we consider a general model that describes how street vendors arrived at their present locations. Specifically, we consider two families of pure birth processes: one with a constant arrival rate (or “birth rate”) in which the Poisson assumption holds and one with a linear arrival rate in which the Poisson assumption does not hold.
[…]
Now suppose instead that larger markets grow faster than smaller markets, an empirical phenomenon known as preferential attachment. Then the time between vendor arrivals may be better approximated by an exponential distribution with a linear arrival rate, or Yule process (Yule 1925). That is, let $N^{t}_{i}(A)$ denote the number of vendors in area $A$ at time $t$ so that the time between vendor arrivals at each market follows an exponential distribution with rate $\lambda_{i}\,N^{t}_{i}(A)$. If we assume further that each market started with one vendor at $t=0$, then the number of vendors at the time the survey was conducted ($t=1$), $N_{i}(A)$, would follow a negative binomial distribution with mean $\mu_{i}=\exp(\lambda_{i})\,M(A)$ and variance $\mu_{i}\bigr{(}\mu_{i}-M(A)\bigr{)}\,/\,M(A)$. i.e.,
$N_{i}(A)\sim\ \text{Negative Binomial}\left(\mu_{i},\ \mu_{i}\,\frac{\mu_{i}-M(A)}{M(A)}\right)$
where we parameterize the negative binomial distribution by its mean and variance. In this case, $N_{i}(A)$ exhibits extra-Poisson variation, and the results of Section 3.2 do not hold. See Ross (2014, chap. 6) for details. [^9] Note that we define the negative binomial and negative hypergeometric distributions as the total number of trials required to get a predetermined number of successes. The negative binomial assumes sampling with replacement, while the negative hypergeometric distribution assumes sampling without replacement.
To derive the standard errors, suppose a portion of the markets, $p$, were selected for the survey so that
$n_{i}(A)\sim\ \text{Negative Binomial}\left(p\,\mu_{i},\ p\,\mu_{i}\,\frac{\mu_{i}-M(A)}{M(A)}\right)$
The conditional distribution of $n_{1}(A)$ given $N_{1}(A)$ is then
$n_{1}(A)\ |\ N_{1}(A)\sim\ \text{Negative Hypergeometric}\bigl{(}N_{1}(A)-1,\ M(A)-1,\ p\,M(A)\bigl{)}$
with conditional mean and variance
$\mathbb{E}\left[n_{1}(A)\ |\ N_{1}(A)\right]=p\,N_{1}(A)\qquad\text{and}\qquad\mathbb{V}\bigr{(}n_{1}(A)\ |\ N_{1}(A)\bigr{)}=w_{1}\,p\,(1-p)\,N_{1}(A)$
so that under increasing domain asymptotics, the standard error of
$\hat{\theta}=\frac{N_{1}(A)\ n_{0}(A)}{n_{1}(A)}$
is asymptotically normal with mean $\theta=\mu_{0}$ and standard error
$\text{SE}(\hat{\theta})=\theta\sqrt{w_{0}\left(\frac{1}{p\,\theta}\right)+w_{1}\left(\frac{1-p}{p\,N_{1}(A)}\right)}$
where
$w_{0}=\dfrac{\theta-M(A)}{M(A)}\qquad\text{and}\qquad w_{1}=\dfrac{N_{1}(A)-M(A)}{M(A)+1}$
See Section 7.3 for details.
Note that this standard error is identical to the standard error from Section 3.2 except for the weights $w_{0}$ and $w_{1}$. These weights are the result of the extra-Poisson variation induced by preferential attachment—i.e., the linear “birth” rate in which the arrival of new vendors is proportional to the number of vendors in the market.
The weights $w_{0}$ and $w_{1}$ correspond to the relative number of vendors without permits and with permits per market respectively. These weights cannot be estimated from the survey data alone. However, we believe that $\theta$ and $N_{1}(A)$ are no more than 10 times larger than $M(A)$. It follows that $w_{0}$ and $w_{1}$ are less than 9, and thus the standard errors from Section 3.2 and the margins of error from Table 1 are at most be 3 times larger if the Poisson assumption is violated and the Yule process holds. We suspect this bound is conservative, however, and the typical market only contains one or two vendors.
5. Discussion
We conclude by discussing the implications of our estimates. In particular, we assess the completeness of the New York City Office of Administrative Trials and Hearings (OATH) and American Community Survey (ACS) data. Our estimates suggest that while both data sets are spatially representative, they miss the majority of street vendors. We provide several explanations for this discrepancy.
We estimate 23,000 vendors work in New York City. In comparison, the 2017-2021 ACS estimates 4,634 people work the occupation of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers in New York City. This suggests that according to the ACS, there are at most 4,634 street vendors in New York City, and thus our finding indicate that the ACS estimate misses 80% of street vendors or more.
A fairer comparison might account for the fact that the standard error of our estimate is approximately 1,000, and the standard error of the ACS estimate is approximately 488, which we obtained from the replicate weights. Using the lower limit of a 95% confidence interval for our estimate (19,000) and the upper 95% confidence limit for the ACS estimate (5,610), our findings suggest the ACS misses roughly two thirds of street vendors or more.
Note the importance of the assumptions stated in Section 3.2 since other derivations of ratio estimation may imply smaller or larger standard errors, from which we might conclude the ACS misses a smaller or larger portion of the street vending population. For example, if the standard errors from Section 4.2 hold, then the confidence intervals would overlap if there were more than 225 vendors in the average market. We believe this is highly unlikely, suggesting that even if the Poisson assumption is violated, the two estimates are inconsistent.
Additional evidence comes from the OATH data. The OATH data indicate that approximately 2,500 (unique) vendors are summoned to court for allegedly violating laws related to vending each year. If the ACS estimate of at most 4,634 vendors holds, then more than half of all vendors are summoned to court each year. If our estimate of 23,000 vendors holds, then approximately one in ten vendors are summoned to court each year. We believe the latter is more realistic.
There may be several reasons why our estimates are significantly larger than the ACS. The fact that the Street Vendor Project distributed relief aid may have encouraged respondents who are unlikely to respond to the ACS. Moreover, the Street Vendor Project, through membership lists, referrals, and canvassing may have covered individuals missed by the Census Bureau’s sampling frame.
Another explanation is that the Street Vendor Project survey and the ACS may be measuring different populations. For example, the ACS estimate is the average number of vendors between 2017 and 2021, while our estimate reflects 2020 and 2021. For another example, the Census Bureau determines the occupation of an ACS respondent by autocoding responses to write-in questions. It is possible that the Street Vendor Project’s definition of a vendor is more inclusive. Indeed, street vending often provides supplemental income, and thus the discrepancy may reflect the broader challenge of studying the gig economy. In this case, either estimate may be relevant, depending on the intended use case.

### v4

@@
 Second, we describe a survey that was administered by the Street Vendor Project while distributing Coronavirus relief aid to vendors operating in New York City both with and without a license or permit.
-Third, we calculate the total number of vendors using ratio estimation, finding approximately 23,000 street vendors operate in New York City—20,500 mobile food vendors and 2,300 general merchandise vendors—with one third located in just six ZIP Codes—11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan.
-Fourth, we provide a theoretical justification of our estimates based on the theory of point processes.
+Third, we review ratio estimation and provide a theoretical justification based on the theory of point processes.
+Fourth, we use ratio estimation to calculate the total number of vendors, finding approximately 23,000 street vendors operate in New York City—20,500 mobile food vendors and 2,400 general merchandise vendors—with one third located in just six ZIP Codes—11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan.
 Finally, we evaluate the accuracy of the ratio estimator when the distribution of vendors is explained by a Poisson or Yule process, and we discuss several policy implications.
@@
 Fines paid by vendors provide approximately \$200,000 each year according to data from the New York City Office of Administrative Trials and Hearings.
-Perhaps most importantly, street vending historically benefits underserved communities, both because vendors operate in neighborhoods with limited access to traditional stores and because vending is one of a handful of occupations in which New Yorkers of all backgrounds, immigrants in particular, are able to achieve economic mobility and a chance at the American dream (Burrows and Wallace 1998, chap.
+Perhaps most importantly, street vending historically benefits underserved communities, both because vendors operate in neighborhoods with limited access to traditional stores and because vending is one of a handful of occupations in which New Yorkers of all backgrounds, immigrants in particular, can achieve economic mobility and a chance at the American dream (Burrows and Wallace 1998, chap.
 42).
@@
 We present our work in four sections.
-In Section 2, we review the process by which vendors receive permits and licenses, and we describe a survey administered by the Street Vendor Project at the Urban Justice Center while distributing Coronavirus relief aid to vendors operating both with and without a license or permit.
+In Section 2, we review the process by which vendors receive licenses and permits, and we describe a survey administered by the Street Vendor Project at the Urban Justice Center while distributing Coronavirus relief aid to vendors operating both with and without a license or permit.
 In Section 3, we use the fact that the number of licenses and permits are limited by law to construct a ratio estimator for the number of street vendors.
-We find approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,300 general merchandise vendors.
+We then provide a theoretical justification of our estimates, assuming the spatial distribution of survey respondents is representative and well-approximated by a set of inhomogeneous Poisson processes.
+We find approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,400 general merchandise vendors.
 One third are located in just six ZIP Codes: 11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan.
-We then provide a theoretical justification of our estimates, assuming the spatial distribution of survey respondents is representative and well-approximated by a family of inhomogeneous Poisson processes.
-In Section 4, we examine the validity of these assumptions.
+We also derive standard errors.
+In Section 4, we examine the validity of our assumptions.
 We evaluate whether the spatial distribution of respondents is representative by comparing it to the American Community Survey and administrative data, and we evaluate the Poisson process assumption by considering a more general model in which the arrival of street vendors follows a pure birth process, of which the Poisson and Yule processes are special cases.
+The Yule process may be appropriate if street vendors cluster within markets, in which case the data might have extra-Poisson variation.
+We rederive the standard errors from Section 3 to account for this variation.
 In Section 5, we discuss some implications of our estimates.
@@
 Our estimates suggest the American Community Survey misses the majority of New York City street vendors.
-2. We consider two types of vendors: food vendors and merchandise vendors
+2. Background and Survey Design
 A street vendor is any individual who sells goods from a mobile vending unit instead of a store.
@@
 We refer to these vendors as food vendors and merchandise vendors, respectively.
-In Section 2.1, we review the process by which vendors receive permits and licenses as required by law.
+In Section 2.1, we review the process by which vendors receive licenses and permits as required by law.
 In Section 2.2, we describe a survey administered by the Street Vendor Project while distributing Coronavirus relief aid to vendors operating in New York City both with and without a license or permit.
-2.1 The number of licenses and permits are limited by local law
+2.1 Background
 A long list of rules determines the individuals, locations, and time periods during which vendors can legally sell their goods in New York City.
@@
 Constitution so that the number of First Amendment vendors cannot be restricted by law.
-The final exception are concessionaires that operate on New York City parkland through an NYC Parks permit or license.
+The final exception are concessionaires that operate on New York City parkland through an NYC Parks license or permit.
 According to the New York City Department of Parks and Recreation, there are approximately 400 concessions within New York City parks.
@@
 Our estimates are limited to food and merchandise vendors operating outside of parks.
-2.2 The Street Vendor Project surveyed two thousand street vendors in 2021
+2.2 Survey Design
 The Street Vendor Project (SVP), part of the Urban Justice Center, is a non-profit organization that advocates on behalf of New York City street vendors.
@@
 Survey operations continued until \$2,415,000 in funds were distributed, yielding 2,060 responses.
-The survey included 100 questions and was conducted in eight different languages: Arabic, Bangla, Cantonese, English, French, Mandarin, Spanish, Tibetan, and Wolof.
+The survey included 100 questions and was conducted in nine different languages: Arabic, Bangla, Cantonese, English, French, Mandarin, Spanish, Tibetan, and Wolof.
 [^4] Vendors who spoke other languages, such as Dari, Farsi, Russian, and Turkish, typically spoke English with a high level of fluency, and those surveys were conducted in English.
-Similarly, vendors who spoke indigenous languages from Latin America, such as Nahuatl or Quechua, often spoke Spanish with a high level of fluency.
+Similarly, vendors who spoke indigenous languages from Latin America, such as Nahuatl and Quechua, typically spoke Spanish with a high level of fluency.
 The survey items solicited a variety of information from vendors, such as logistical information (e.g., vending location, residential location, and frequency of operation), economic information (e.g., items sold, income, and expenses), and demographic information (e.g., age, race, ethnicity, and immigration status).
@@
 Approximately five percent of vendors did not identify their vending location.
-These vendors are excluded from Table 1 and Figure 1, although doing so does not meaningfully change the estimated number of vendors that operate in New York City.
-Veteran merchandise vendors are also excluded from Table 1 and Figure 1, although these vendors are reflected in our city-wide estimate of 23,000 vendors.
-3. We calculate the number of street vendors using ratio estimation
-We estimate approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,300 general merchandise vendors.
+These vendors are included in the New York City total, but they are excluded from the neighborhood estimates in Table 1 and Figure 1.
+(For this reason, the Respondent and Population columns do not sum exactly to the New York City total.) Veteran merchandise vendors are excluded entirely from Table 1 and Figure 1, although these vendors are reflected in our overall estimate of 23,000 vendors.
+3. Methodology and Data Analysis
+We estimate approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,400 general merchandise vendors.
 We arrive at the estimate by using ratio estimation, which leverages the fact that the number of licenses and permits are limited by local law.
@@
 Inference is conducted using increasing domain asymptotics, and the details are outlined in Section 7.3.
-3.1 We estimate 23,000 street vendors operate in New York City
+In Section 3.3, we apply the methodology to the survey data discussed in Section 2.2.
+3.1 Ratio Estimation
 Ratio estimation is a common approach for estimating the size of a population.
@@
 4) for an introduction.
-The popularity of ratio estimation is due in large part to its simplicity: Among the $1,400$ food vendors surveyed, roughly a quarter had a permit.
-Suppose this ratio is representative in that one quarter of all food vendors have a permit.
-It follows that the total number of food vendors is $20,500\approx 5,100/(1/4)$ because one quarter of $20,500$ is approximately $5,100$, the number of permits available.
-By similar reasoning—i.e., exploiting the fact that the number of general merchandise licenses is limited for non-veterans—the number of non-veteran merchandise vendors is estimated to be approximately $1,300$.
-When combined with the estimated $1,000$ veteran merchandise vendors (see Section 2.1) and $20,500$ food vendors, the total number of street vendors is approximately $23,000$.
-The main assumption of ratio estimation is that the survey data are representative in the sense that, were survey operations to continue, the percentage of respondents with a permit (in the case of food vendors) or a license (in the case of merchandise vendors) would remain the same.
-If we further assume that the spatial distribution of respondents is representative, we can also estimate the total number of vendors by neighborhood.
-[^5] We use the same ratio for each neighborhood.
-Ratio estimation by neighborhood is not possible since only the locations of the vendors who participated in the survey are known.
-These estimates are listed in Section 7.1, Table 1.
-We find that approximately one quarter of street vendors are located in the Chelsea and Clinton neighborhoods of Manhattan.
-Another quarter is located in West Queens and Lower Manhattan.
-We check our estimates by comparing them to a second, independent assessment conducted by the Street Vendor Project in the Bronx on May 13 and 15, 2022, in which the number of visible street vendors was documented through on-the-ground observations.
-The independent assessment identified 188 total street vendors in the Fordham area (Fordham Road BID and Street Vendor Project 2024), whereas Table 1 indicates 206 vendors in Bronx Park and Fordham—an estimate approximately ten percent higher.
-To demonstrate that this difference can be explained by sampling variation, we derive the standard error of the ratio estimate in Section 3.2.
-3.2 We model the location of street vendors as a Poisson process
-We provide a point process justification of the ratio estimation procedure described in Section 3.1. The purpose is to examine the underlying assumptions and derive the standard error of the estimate.
-Let $\{\Pi_{i}\}$ denote a family of inhomogeneous spatial Poisson processes, referencing the location $x\in A\subset\mathbb{R}^{2}$ of each street vendor at the time the survey was conducted.
-The index $i$ denotes the status of the vendor.
-To fix ideas, let $i=0$ indicate mobile food vendors with a permit, and let $i=1$ indicate mobile food vendors without a permit.
+We provide a simple explanation based on cross-multiplication, also called the rule of three.
+The purpose is to highlight the main assumption.
+A more rigorous derivation is provided in Section 3.2.
+Consider a fixed region $A$, and let $N_{i}(A)$ denote the number of individuals in region $A$ of type $i$.
+To fix ideas, let $A$ be New York City, let $i=0$ indicate mobile food vendors with a permit, and let $i=1$ indicate mobile food vendors without a permit.
+Also, let $n_{i}(A)$ denote the number of respondents of type $i$.
+We observe $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(A)$ from which we estimate $\theta(A)=\mathbb{E}\left[N_{0}(A)\right]$, the expected number of vendors without a permit.
+The expected number of mobile food vendors with or without a permit is then $\tau(A)=\theta(A)+N_{1}(A)$.
+The ratio estimator of $\theta(A)$ is
+$\hat{\theta}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}$
+It is called the ratio estimator because it depends on the ratio of random variables $n_{0}(A)\,/\,n_{1}(A)$.
+(The total number of permitted vendors, $N_{1}(A)$, is considered fixed.) The estimator is accurate when the response rate, $p$, is approximately the same for each type, i.e.,
+$p\approx\frac{n_{0}(A)}{N_{0}(A)}\approx\frac{n_{1}(A)}{N_{1}(A)}$
+in which case $\hat{\theta}(A)$ is the solution for $N_{0}(A)$ in terms of $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(A)$, obtained by cross-multiplication.
+The main assumption of ratio estimation is that the survey data are representative in the sense that the response rate does not depend on whether a mobile food vendor has a permit.
+If we further assume that the response rate does not vary by subregion, we can also estimate the total number of mobile food vendors by subregion.
+For example, suppose the response rate in subregion $B\subset A$ is also approximately equal to $p$, e.g.,
+$p\approx\frac{n_{1}(A)}{N_{1}(A)}\approx\frac{n_{0}(B)}{N_{0}(B)}$
+Then we can estimate $\theta(B)$ by the subregion estimator
+$\tilde{\theta}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}$
+where $\tilde{\theta}(B)$ is the solution for $N_{0}(B)$ in terms of $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(B)$.
+The subregion estimator is necessary when $N_{1}(B)$ is not observed and $\hat{\theta}(B)$ cannot be calculated directly.
+This is the case with the survey data described in Section 2.2.
+We observe $N_{1}(A)$, the number of permits city-wide (i.e., within region $A$), but not $N_{1}(B)$, the number of permits within subregion $B$.
+(We also observe $n_{0}(B)$ and $n_{1}(B)$.)
+The subregion estimator $\tilde{\theta}(B)$ works by replacing the term $N_{1}(B)$ in the ratio estimator $\hat{\theta}(B)$ with the estimate
+$\tilde{N}_{1}(B)=\dfrac{N_{1}(A)\,n_{1}(B)}{n_{1}(A)}$
+The estimate $\tilde{N}_{1}(B)$ is also used to estimate the total number of mobile food vendors expected in subregion $B$. That is, we estimate $\tau(B)=\theta(B)+N_{1}(B)$ with
+$\tilde{\tau}(B)=\tilde{\theta}(B)+\tilde{N}_{1}(B)=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(A)}$
+3.2 Model
+We provide a point process justification of the ratio estimation procedure described in Section 3.1.
+The purpose is to more closely examine the underlying assumptions and derive the standard errors of the estimates.
+Let $\Pi_{i}$ denote an inhomogeneous spatial Poisson process referencing the location $x\in A\subset\mathbb{R}^{2}$ of each street vendor with status $i$ at the time the survey was conducted.
+As in Section 3.1, we fix ideas by letting $i=0$ indicate mobile food vendors with a permit and $i=1$ indicate mobile food vendors without a permit.
 The Poisson process assumption may be justified by the law of rare events.
@@
 Whether a vendor is located within a partition has a vanishingly small probability, such that the number of vendors in any area is well approximated by a Poisson distribution.
-This approximation is accurate even if these probabilities are weakly dependent.
-For example, see Freedman (1974), Chen (1975), and Serfling (1975).
-Note that other justifications are possible.
+This approximation is accurate even if these probabilities are weakly dependent, see Freedman (1974).
+Other justifications are possible.
 For example, see Section 4.2 for a justification of the Poisson process based on the theory of birth processes.
@@
 […]
-so that by conditioning on $N_{1}(A)$, we arrive at the following probability model for the number of respondents with and without permits
-$n_{0}(A)\sim$ $\ \text{Poisson}\big{(}p\,\theta\big{)}$
-$n_{1}(A)\ |\ N_{1}(A)\sim$ $\ \text{Binomial}\big{(}p,\,N_{1}(A)\big{)}$
-The maximum likelihood estimates for $p$ and $\theta$ are
-$\hat{p}=\dfrac{n_{1}(A)}{N_{1}(A)}\qquad\text{and}\qquad\hat{\theta}=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}$
-The estimator $\hat{\theta}$ is called the ratio estimator because it depends on the ratio of random variables $n_{0}(A)\,/\,n_{1}(A)$.
-(The total number of permitted vendors, $N_{1}(A)$, is considered fixed.) Using increasing domain asymptotics, the estimator is asymptotically normal with mean $\theta$ and standard error
-$\text{SE}(\hat{\theta})=\theta\sqrt{\dfrac{1}{p\,\theta}+\dfrac{1-p}{p\,N_{1}(A)}}$
-Substituting $\hat{p}$ and $\hat{\theta}$ for $p$ and $\theta$ yields the following plug-in estimator for the standard error,
-$\hat{\text{SE}}(\hat{\theta})=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(A)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
-See Section 7.3 for details.
-The margin of error reported in Section 7.1, Table 1 is two standard errors, half of the width of a 95% confidence interval.
-Note that the margin of error for the 206 vendors estimated in Bronx Park and Fordham is 44, and thus this estimate is consistent with the 188 street vendors counted independently by Fordham Road BID and Street Vendor Project (2024) as described in Section 3.1.
-4. We examine the validity of our assumptions using additional information
+so that by conditioning on $N_{1}(B)$, we arrive at the following probability model for the number of respondents with and without permits
+$n_{0}(B)\sim$ $\ \text{Poisson}\big{(}p\,\theta(B)\big{)}$
+$n_{1}(B)\ |\ N_{1}(B)\sim$ $\ \text{Binomial}\big{(}p,\,N_{1}(B)\big{)}$
+The maximum likelihood estimates for $p$ and $\theta(A)$ are
+$\hat{p}=\dfrac{n_{1}(A)}{N_{1}(A)}\qquad\text{and}\qquad\hat{\theta}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}$
+Using increasing domain asymptotics, the estimator is asymptotically normal with mean $\theta(A)$ and standard error
+$\text{SE}[\hat{\theta}(A)]=\theta(A)\sqrt{\dfrac{1}{p\,\theta(A)}+\dfrac{1-p}{p\,N_{1}(A)}}$
+Substituting $\hat{p}$ and $\hat{\theta}(A)$ for $p$ and $\theta(A)$ yields the following plug-in estimator for the standard error,
+$\hat{\text{SE}}[\hat{\theta}(A)]=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(A)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
+See Section 7.3.2 for details.
+The maximum likelihood estimate for $\theta(B)$ is
+$\tilde{\theta}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}$
+which is asymptotically normal with mean $\theta(B)$ and standard error
+$\text{SE}[\tilde{\theta}(B)]=\theta(B)\sqrt{\frac{1}{p\,\theta(B)}+\frac{1-p}{p\,N_{1}(A)}}$
+The plug-in estimator for the standard error is
+$\tilde{\text{SE}}[\tilde{\theta}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
+To determine the subregion total $\tau(B)=\theta(B)+N_{1}(B)$, we estimate $\theta(B)$ and $N_{1}(B)$ with $\tilde{\theta}(B)$ and $\tilde{N}_{1}(B)=N_{1}(A)\,n_{1}(B)\,/\,n_{1}(A)$, yielding
+$\tilde{\tau}(B)=\tilde{\theta}(B)+\tilde{N}_{1}(B)=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(A)}=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(B)+n_{1}(A\backslash B)}$
+where $A\backslash B$ denotes the compliment of subregion $B$.
+The asymptotic standard error is
+$\text{SE}[\tilde{\tau}(B)]=\theta(B)\sqrt{\frac{1}{p\,\theta(B)}+\frac{1-p}{p\,N_{1}(A)}\left(1+\frac{N_{1}(B)\,N_{1}(A\backslash B)}{\theta(B)^{2}}\right)}$
+and the plug-in estimator is
+$\tilde{\text{SE}}[\tilde{\tau}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\left(\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}\right)\left(1+\dfrac{n_{1}(B)\,n_{1}(A\backslash B)}{n_{0}(B)^{2}}\right)}$
+See Section 7.3.3 for details.
+3.3 Data Analysis
+We use the methodology described in Section 3.1 and 3.2 to estimate the number of street vendors in New York City.
+According to the survey data described in Section 2.2, $n_{1}(A)=349$ mobile food vendor respondents indicated they had a permit, and $n_{0}(A)=1,051$ indicated they did not.
+Since there are $N_{1}(A)=5,100$ permits available, it follows that the number of vendors without permits is 15,3505,100×1,051/ 349formulae-sequence153505100105134915,350$\approx$ 5,100$\times$ 1,051\,/\,34915 , 350 ≈ 5 , 100 × 1 , 051 / 349.
+Combined with the $5,100$ permitted vendors, the total number of mobile food vendors is approximately $20,500$.
+By similar reasoning—i.e., exploiting the fact that the number of general merchandise licenses is limited for non-veterans—the number of non-veteran merchandise vendors is estimated to be approximately $1,400$.
+When combined with the estimated $1,000$ veteran merchandise vendors (see Section 2.1) and $20,500$ food vendors, the total number of street vendors is approximately $23,000$.
+Estimates for select subregions are listed in Section 7.1, Table 1.
+We find that approximately one quarter of street vendors are located in the Manhattan neighborhoods of Chelsea, Clinton, and Lower Manhattan.
+Another quarter is located in West and North Queens.
+We also report the margin of error (i.e., two standard errors, half of the width of a 95% confidence interval).
+We check these estimates by comparing them to a second, independent assessment conducted by the Street Vendor Project in the Bronx on May 13 and 15, 2022, in which the number of street vendors was documented through on-the-ground observations.
+The independent assessment identified 188 total street vendors near Fordham Road (Fordham Road BID and Street Vendor Project 2024).
+According to the survey data, 17 respondents indicated that they operate near Fordham Road, yielding an estimate of 152 street vendors total with a standard error of 54.
+We conclude this estimate is consistent with the 188 street vendors counted independently by Fordham Road BID and Street Vendor Project (2024).
+4. Model Validation
 We make two assumptions in Section 3.
 The main assumption is that the status and spatial distribution of the respondents is representative.
-This assumption is necessary to ensure the ratio estimator is consistent.
+This assumption is necessary to ensure the ratio estimator is accurate in large samples (i.e., consistent).
 A secondary assumption is that the spatial distribution of vendors is well described by a family of inhomogeneous spatial Poisson processes.
@@
 Since neither can be evaluated from the survey data alone, we consider additional information.
-In Section 4.1, we compare the spatial distribution of the respondents to data from the American Community Survey Public Use Micro Sample 2017-2021 and the New York City Office of Administrative Trials and Hearings.
+In Section 4.1, we compare the spatial distribution of the respondents to data from the New York City Office of Administrative Trials and Hearings and the American Community Survey.
 In Section 4.2, we model the arrival of street vendors as a pure birth process, and we derive a more general formula for the standard errors in which the amount of extra-Poisson variation can be determined by the number of vendors that cluster within markets.
@@
 The advantage of the birth process approach taken in Section 4.2 is that it provides a plausible explanation for why the Poisson assumption made in Section 3.2 may not hold in practice.
-4.1 The Street Vendor Project survey is spatially representative
+4.1 Representativeness
 The main assumption in Section 3 is that the status and spatial distribution of the respondents is representative (i.e., the response probability is spatially uncorrelated with the mean measure).
@@
 See Daley and Vere-Jones (2007) for a general statement of Campbell’s Theorem.
-If this assumption is violated, then $\hat{\theta}$ may be inconsistent.
-For example, if respondents are more likely to come from locations in which a higher proportion of vendors have permits, then the ratio $n_{0}(A)\,/\,n_{1}(A)$ would not be representative, and $\hat{\theta}$ would likely underestimate $\theta$ even in large samples.
-We think this assumption is reasonable because the Street Vendor Project canvassed New York City to distribute a large amount of relief aid.
+If this assumption is violated, then $\hat{\theta}(A)$ may be inconsistent.
+For example, if respondents are more likely to come from locations in which a higher proportion of vendors have permits, then the ratio $n_{0}(A)\,/\,n_{1}(A)$ would not be representative, and $\hat{\theta}(A)$ would likely underestimate $\theta(A)$ even in large samples.
+We think the representativeness assumption is reasonable because the Street Vendor Project canvassed New York City to distribute a large amount of relief aid.
 We argue that trust in SVP and its mission, along with the aid, removed many barriers that typically prevent survey operations from reaching hard-to-reach populations.
@@
 We thus compare the spatial distribution of vending locations reported by respondents to the spatial distribution of vendors who allegedly violated street vending laws during the year 2021.
-We consider the location of an individual’s first violation in 2021, reported in administrative data publicly available from the New York City Office of Administrative Trials and Hearings (OATH).
-We also compare the residencies reported by the respondents to the residencies of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers (Standard Occupational Classification 41-9091) estimated in the five year American Community Survey Public Use Micro Sample, 2017-2021 (ACS).
+We consider the location of an individual’s first violation in 2021, reported in administrative data from the NYC Office of Administrative Trials and Hearings (OATH) (accessed 2024-04-03).
+We also compare the residencies reported by the respondents to the residencies of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers (Standard Occupational Classification 41-9091) estimated in the 2017-2021 American Community Survey (ACS) Public Use Micro Sample from the U.S.
+Census Bureau (accessed 2022-12-22).
 These comparisons are visualized in Section 7.2, Figure 1.
@@
 We also note that even if this discrepancy suggests a location was missed by the Street Vendor Project, our main assumption is not necessarily violated.
-That is because this undercount may be compensated by an overcount in other areas (e.g., the Upper East Side of Manhattan or East Brooklyn) such that our overall estimate is still accurate.
-4.2 We model the arrival of street vendors as a pure birth process
+That is because the undercount in some areas may be compensated by an overcount in others (e.g., the Upper East Side of Manhattan or East Brooklyn) such that our overall estimate is still accurate.
+4.2 Overdispersion
 The second assumption in Section 3.2 is that the spatial distribution of vendors is well approximated by a family of inhomogeneous Poisson processes.
-We consider this assumption secondary because the consistency of $\hat{\theta}$ holds under the wide class of point processes described in Section 4.1.
+We consider this assumption secondary because the consistency of $\hat{\theta}(A)$ holds under a wide class of point processes described in Section 4.1.
 That said, the standard errors derived in Section 3.2 do not necessarily hold when the Poisson assumption is violated.
 [^6] The ratio estimator may no longer be asymptotically efficient as well.
-i.e., In Section 7.3, we show that under the Poisson process assumption, the ratio estimator maximizes the likelihood, which implies the ratio estimator is asymptotically efficient.
+In Section 7.3, we show that under the Poisson process assumption, the ratio estimator maximizes the likelihood, which implies it is asymptotically efficient.
 However, the ratio estimator may not maximize the likelihood when the Poisson process assumption is violated.
@@
 Then the time between vendor arrivals may be better approximated by an exponential distribution with a linear arrival rate, or Yule process (Yule 1925).
-That is, let $N^{t}_{i}(A)$ denote the number of vendors in area $A$ at time $t$ so that the time between vendor arrivals at each market follows an exponential distribution with rate $\lambda_{i}\,N^{t}_{i}(A)$.
-If we assume further that each market started with one vendor at $t=0$, then the number of vendors at the time the survey was conducted ($t=1$), $N_{i}(A)$, would follow a negative binomial distribution with mean $\mu_{i}=\exp(\lambda_{i})\,M(A)$ and variance $\mu_{i}\bigr{(}\mu_{i}-M(A)\bigr{)}\,/\,M(A)$.
+That is, let $N^{t}_{i}(B)$ denote the number of vendors in subregion $B$ at time $t$ so that the time between vendor arrivals at each market follows an exponential distribution with rate $\lambda_{i}\,N^{t}_{i}(B)$.
+If we assume further that each market started with one vendor at $t=0$, then the number of vendors at the time the survey was conducted ($t=1$), $N_{i}(B)$, would follow a negative binomial distribution with mean $\mu_{i}(B)=\exp(\lambda_{i})\,M(B)$ and variance $\mu_{i}(B)\bigr{(}\mu_{i}(B)-M(B)\bigr{)}\,/\,M(B)$.
 i.e.,
-$N_{i}(A)\sim\ \text{Negative Binomial}\left(\mu_{i},\ \mu_{i}\,\frac{\mu_{i}-M(A)}{M(A)}\right)$
+$N_{i}(B)\sim\ \text{Negative Binomial}\left(\mu_{i}(B),\ \mu_{i}(B)\,\frac{\mu_{i}(B)-M(B)}{M(B)}\right)$
 where we parameterize the negative binomial distribution by its mean and variance.
-In this case, $N_{i}(A)$ exhibits extra-Poisson variation, and the results of Section 3.2 do not hold.
+In this case, $N_{i}(B)$ exhibits extra-Poisson variation, and the results of Section 3.2 do not hold.
 See Ross (2014, chap.
@@
 To derive the standard errors, suppose a portion of the markets, $p$, were selected for the survey so that
-$n_{i}(A)\sim\ \text{Negative Binomial}\left(p\,\mu_{i},\ p\,\mu_{i}\,\frac{\mu_{i}-M(A)}{M(A)}\right)$
-The conditional distribution of $n_{1}(A)$ given $N_{1}(A)$ is then
-$n_{1}(A)\ |\ N_{1}(A)\sim\ \text{Negative Hypergeometric}\bigl{(}N_{1}(A)-1,\ M(A)-1,\ p\,M(A)\bigl{)}$
+$n_{i}(B)\sim\ \text{Negative Binomial}\left(p\,\mu_{i}(B),\ p\,\mu_{i}(B)\,\frac{\mu_{i}(B)-M(B)}{M(B)}\right)$
+The conditional distribution of $n_{1}(B)$ given $N_{1}(B)$ is then
+$n_{1}(B)\ |\ N_{1}(B)\sim\ \text{Negative Hypergeometric}\bigl{(}N_{1}(B)-1,\ M(B)-1,\ p\,M(B)\bigl{)}$
 with conditional mean and variance
-$\mathbb{E}\left[n_{1}(A)\ |\ N_{1}(A)\right]=p\,N_{1}(A)\qquad\text{and}\qquad\mathbb{V}\bigr{(}n_{1}(A)\ |\ N_{1}(A)\bigr{)}=w_{1}\,p\,(1-p)\,N_{1}(A)$
-so that under increasing domain asymptotics, the standard error of
-$\hat{\theta}=\frac{N_{1}(A)\ n_{0}(A)}{n_{1}(A)}$
-is asymptotically normal with mean $\theta=\mu_{0}$ and standard error
-$\text{SE}(\hat{\theta})=\theta\sqrt{w_{0}\left(\frac{1}{p\,\theta}\right)+w_{1}\left(\frac{1-p}{p\,N_{1}(A)}\right)}$
+$\mathbb{E}\left[n_{1}(B)\ |\ N_{1}(B)\right]=p\,N_{1}(B)\qquad\text{and}\qquad\mathbb{V}\bigr{(}n_{1}(B)\ |\ N_{1}(B)\bigr{)}=w_{1}\,p\,(1-p)\,N_{1}(B)$
+so that under increasing domain asymptotics, the ratio estimator
+$\hat{\theta}(A)=\frac{N_{1}(A)\ n_{0}(A)}{n_{1}(A)}$
+is asymptotically normal with mean $\theta(A)=\mu_{0}(A)$ and standard error
+$\text{SE}[\hat{\theta}(A)]=\theta(A)\sqrt{w_{0}\left(\frac{1}{p\,\theta(A)}\right)+w_{1}\left(\frac{1-p}{p\,N_{1}(A)}\right)}$
 where
-$w_{0}=\dfrac{\theta-M(A)}{M(A)}\qquad\text{and}\qquad w_{1}=\dfrac{N_{1}(A)-M(A)}{M(A)+1}$
-See Section 7.3 for details.
-Note that this standard error is identical to the standard error from Section 3.2 except for the weights $w_{0}$ and $w_{1}$.
-These weights are the result of the extra-Poisson variation induced by preferential attachment—i.e., the linear “birth” rate in which the arrival of new vendors is proportional to the number of vendors in the market.
+$w_{0}=\dfrac{\theta(A)-M(A)}{M(A)}\qquad\text{and}\qquad w_{1}=\dfrac{N_{1}(A)-M(A)}{M(A)+1}$
+See Section 7.3.4 for details. Note that the weights $w_{0}$ and $w_{1}$ depend on $A$, although this dependence is not reflected in the notation.
+The standard error derived in this section for $\hat{\theta}(A)$ is identical to the standard error for $\hat{\theta}(A)$ from Section 3.2 except for the weights $w_{0}$ and $w_{1}$.
+These weights are the result of the extra-Poisson variation induced by preferential attachment—that is, the linear “birth” rate in which the arrival of new vendors is proportional to the number of vendors in the market.
 The weights $w_{0}$ and $w_{1}$ correspond to the relative number of vendors without permits and with permits per market respectively.
 These weights cannot be estimated from the survey data alone.
-However, we believe that $\theta$ and $N_{1}(A)$ are no more than 10 times larger than $M(A)$.
-It follows that $w_{0}$ and $w_{1}$ are less than 9, and thus the standard errors from Section 3.2 and the margins of error from Table 1 are at most be 3 times larger if the Poisson assumption is violated and the Yule process holds.
-We suspect this bound is conservative, however, and the typical market only contains one or two vendors.
+However, we believe that $\theta(A)$ and $N_{1}(A)$ are no more than 10 times larger than $M(A)$.
+It follows that $w_{0}$ and $w_{1}$ are less than 9, and thus the standard errors from Section 3.2 and the margins of error from Table 1 are at most 3 times larger if the Poisson assumption is violated and the Yule process holds.
+We suspect this bound is conservative, however, and the typical market contains only one or two vendors.
+The same behavior is exhibited by the subregion estimator of $\theta(B)$,
+$\tilde{\theta}(B)=\dfrac{N_{1}(A)\ n_{0}(B)}{n_{1}(A)}$
+which is asymptotically normal with mean $\theta(B)=\mu_{0}(B)$ and standard error
+$\text{SE}[\tilde{\theta}(B)]=\theta(B)\sqrt{v_{0}\left(\frac{1}{p\,\theta(B)}\right)+w_{1}\left(\frac{1-p}{p\,N_{1}(A)}\right)}$
+where
+$v_{0}=\dfrac{\theta(B)-M(B)}{M(B)}$
+The estimator for the subregion total, $\tilde{\tau}(B)=\tilde{\theta}(B)+\tilde{N}_{1}(B)$, is also asymptotically normal with mean $\tau(B)$ and standard error
+$\text{SE}[\tilde{\tau}(B)]=\theta(B)\sqrt{v_{0}\left(\frac{1}{p\,\theta(B)}\right)+\left(\frac{v_{1}\,N_{1}(B)\left(N_{1}(A\backslash B)-\theta(B)\right)^{2}+v_{2}\,N_{1}(A\backslash B)\left(N_{1}(B)+\theta(B)\right)^{2}}{N_{1}(A)\,\theta(B)^{2}}\right)\left(\frac{1-p}{p\,N_{1}(A)}\right)}$
+where
+$v_{1}=\dfrac{N_{1}(B)-M(B)}{M(B)+1}\qquad\text{and}\qquad v_{2}=\dfrac{N_{1}(A\backslash B)-M(A\backslash B)}{M(A\backslash B)+1}$
+See Section 7.3.4 for details.
 5. Discussion
@@
 We estimate 23,000 vendors work in New York City.
-In comparison, the 2017-2021 ACS estimates 4,634 people work the occupation of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers in New York City.
-This suggests that according to the ACS, there are at most 4,634 street vendors in New York City, and thus our finding indicate that the ACS estimate misses 80% of street vendors or more.
+In comparison, the 2017-2021 ACS estimates 4,634 individuals work the occupation of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers in New York City.
+This suggests that, according to the ACS, there are at most 4,634 street vendors in New York City, and thus our finding indicates that the ACS estimate misses 80% of street vendors or more.
 A fairer comparison might account for the fact that the standard error of our estimate is approximately 1,000, and the standard error of the ACS estimate is approximately 488, which we obtained from the replicate weights.
-Using the lower limit of a 95% confidence interval for our estimate (19,000) and the upper 95% confidence limit for the ACS estimate (5,610), our findings suggest the ACS misses roughly two thirds of street vendors or more.
+Using the lower limit of a 95% confidence interval for our estimate (21,000) and the upper 95% confidence limit for the ACS estimate (5,610), our findings suggest the ACS misses roughly three quarters of street vendors or more.
 Note the importance of the assumptions stated in Section 3.2 since other derivations of ratio estimation may imply smaller or larger standard errors, from which we might conclude the ACS misses a smaller or larger portion of the street vending population.
-For example, if the standard errors from Section 4.2 hold, then the confidence intervals would overlap if there were more than 225 vendors in the average market.
-We believe this is highly unlikely, suggesting that even if the Poisson assumption is violated, the two estimates are inconsistent.
-Additional evidence comes from the OATH data.
-The OATH data indicate that approximately 2,500 (unique) vendors are summoned to court for allegedly violating laws related to vending each year.
-If the ACS estimate of at most 4,634 vendors holds, then more than half of all vendors are summoned to court each year.
-If our estimate of 23,000 vendors holds, then approximately one in ten vendors are summoned to court each year.
+For example, if the standard errors from Section 4.2 hold, then the confidence intervals would overlap if there were more than 300 vendors in the average market.
+We believe this is highly unlikely, suggesting the two estimates are inconsistent even under overdispersion.
+Additional evidence comes from the OATH data, which indicate approximately 2,500 (unique) vendors are summoned to court each year for violating vending laws.
+If the ACS estimate holds, then more than half of vendors are summoned to court each year.
+If our estimate holds, then approximately one in ten vendors are summoned to court each year.
 We believe the latter is more realistic.
@@
 Moreover, the Street Vendor Project, through membership lists, referrals, and canvassing may have covered individuals missed by the Census Bureau’s sampling frame.
-Another explanation is that the Street Vendor Project survey and the ACS may be measuring different populations.
-For example, the ACS estimate is the average number of vendors between 2017 and 2021, while our estimate reflects 2020 and 2021.
+Another explanation is that the Street Vendor Project survey and the ACS may be measuring different populations or concepts.
+For example, the ACS estimate reflects the average number of vendors between 2017 and 2021, while our estimate reflects 2020 and 2021.
 For another example, the Census Bureau determines the occupation of an ACS respondent by autocoding responses to write-in questions.
 It is possible that the Street Vendor Project’s definition of a vendor is more inclusive.
-Indeed, street vending often provides supplemental income, and thus the discrepancy may reflect the broader challenge of studying the gig economy.
+Street vending often provides supplemental income, and thus the discrepancy may reflect the broader challenge of studying the gig economy.
 In this case, either estimate may be relevant, depending on the intended use case.

### v5

@@
 We estimate the number of street vendors in New York City.
-First, we summarize the process by which vendors receive licenses and permits to operate legally in New York City.
-Second, we describe a survey that was administered by the Street Vendor Project while distributing Coronavirus relief aid to vendors operating in New York City both with and without a license or permit.
-Third, we review ratio estimation and provide a theoretical justification based on the theory of point processes.
-Fourth, we use ratio estimation to calculate the total number of vendors, finding approximately 23,000 street vendors operate in New York City—20,500 mobile food vendors and 2,400 general merchandise vendors—with one third located in just six ZIP Codes—11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan.
-Finally, we evaluate the accuracy of the ratio estimator when the distribution of vendors is explained by a Poisson or Yule process, and we discuss several policy implications.
-In particular, our estimates suggest the American Community Survey misses the majority of New York City street vendors.
+We first summarize the process by which vendors receive licenses and permits to legally operate in New York City.
+We then describe a survey that was administered by the Street Vendor Project while distributing Coronavirus relief aid to vendors operating in New York City both with and without a license or permit.
+Finally, we calculate the total number of vendors using ratio estimation.
+We find that approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,300 general merchandise vendors.
+One third are located in just six ZIP Codes: 11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan.
+We also provide a theoretical justification of our estimates based on the theory of point processes and a discussion of their accuracy and implications.
+In particular, our estimates suggest the American Community Survey fails to cover the majority of New York City street vendors.
 1. Introduction
@@
 They are an iconic part of the urban landscape and a thriving sector of the local economy, contributing millions of dollars in government revenue through taxes, fines, and fees.
-[^1] Fees from mobile food vending licenses and permits provide nearly a million dollars in revenue to New York City each year (Mosher and Turnquist 2024).
-Fines paid by vendors provide approximately \$200,000 each year according to data from the New York City Office of Administrative Trials and Hearings.
 Perhaps most importantly, street vending historically benefits underserved communities, both because vendors operate in neighborhoods with limited access to traditional stores and because vending is one of a handful of occupations in which New Yorkers of all backgrounds, immigrants in particular, can achieve economic mobility and a chance at the American dream (Burrows and Wallace 1998, chap.
@@
 Nevertheless, understanding the size and location of New York’s street vending population is crucial for informing policy and advocacy.
-In this paper, we propose an estimate for the number of street vendors in New York City, including those that operate without a license or permit.
-We present our work in four sections.
+In this paper, we propose and justify an estimator for the number of street vendors in New York City, including those that operate without a license or permit. We present our work in six sections:
 In Section 2, we review the process by which vendors receive licenses and permits, and we describe a survey administered by the Street Vendor Project at the Urban Justice Center while distributing Coronavirus relief aid to vendors operating both with and without a license or permit.
 In Section 3, we use the fact that the number of licenses and permits are limited by law to construct a ratio estimator for the number of street vendors.
-We then provide a theoretical justification of our estimates, assuming the spatial distribution of survey respondents is representative and well-approximated by a set of inhomogeneous Poisson processes.
+We then develop a theoretical justification based on the theory of point processes, which assumes the spatial distribution of survey respondents is representative and well-approximated by a set of inhomogeneous Poisson processes.
+Finally, we derive the standard errors.
+In Section 4, we examine the sensitivity of the ratio estimator to deviations from the underlying assumptions.
+We first consider the case in which the spatial distribution of the survey respondents is not representative, and we examine a more general estimator in which sampling weights adjust for differential nonresponse.
+We then consider the case in which the Poisson process assumption is violated, and we examine a more general model in which the arrival of street vendors follows a pure birth process, of which the Poisson and Yule processes are special cases.
+The Yule process may be more appropriate if street vendors cluster within markets.
+In both cases, we rederive the standard errors.
+In Section 5, we estimate the number of street vendors in New York City.
 We find approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,400 general merchandise vendors.
-One third are located in just six ZIP Codes: 11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan.
-We also derive standard errors.
-In Section 4, we examine the validity of our assumptions.
-We evaluate whether the spatial distribution of respondents is representative by comparing it to the American Community Survey and administrative data, and we evaluate the Poisson process assumption by considering a more general model in which the arrival of street vendors follows a pure birth process, of which the Poisson and Yule processes are special cases.
-The Yule process may be appropriate if street vendors cluster within markets, in which case the data might have extra-Poisson variation.
-We rederive the standard errors from Section 3 to account for this variation.
+We also find that this estimate is somewhat insensitive to deviations from our underlying assumptions.
+For example, we find that if the spatial distribution of survey respondents is not representative, the estimated number of street vendors in New York City could be as low as 21,000 and as high as 32,000.
 In Section 6, we discuss some implications of our estimates.
-In particular, we find that though the spatial distribution of respondents matches the results of the American Community Survey, the two disagree on the size of New York City’s street vending population.
+In particular, we find that although the spatial distribution of respondents matches the results of the American Community Survey, the two disagree on the size of New York City’s street vending population.
 Our estimates suggest the American Community Survey misses the majority of New York City street vendors.
+This finding is important because the American Community Survey is commonly used for policy and advocacy, and our estimates shed light on the broader challenges of measuring the “gig economy.”
 2. Background and Survey Design
@@
 There may be multiple vendors associated with any establishment, and multiple establishments may be owned by a single individual or firm.
-[^2] For example, suppose Sally owns the business Sally’s Salads, which consists of two food trucks, each staffed by three employees.
+For example, suppose Sally owns the business Sally’s Salads, which consists of two food trucks, each staffed by three employees.
 Then Sally’s Salads counts as one business, two establishments, and seven vendors.
@@
 A mobile food vending license authorizes an individual to prepare or serve food from a permitted mobile food vending unit.
-[^3] Note that our estimate reflects the number of street vendors as of 2021, prior to the implementation of Local Law 18 of 2021, which was enacted concurrently.
-Local Law 18 made several changes to the license and permit process.
-These changes do not impact our estimates and therefore are not discussed in this paper.
-There are several types of mobile food vendor permits with varying restrictions.
+There are several versions of a mobile food vendor permit with varying restrictions.
 Of the 5,100 permits available, 200 are borough permits that limit vendors to one of the five boroughs; 100 are reserved for veterans or vendors with a disability; 1,000 are seasonal and valid only from April to October; 1,000 are green cart permits that limit vendors to selling fruit, vegetables, plain nuts, and water; and 2,800 are unrestricted.
@@
 SVP found aid-eligible individuals through membership lists, referrals, and canvassing operations in which SVP affiliates visited neighborhoods.
-Survey operations continued until \$2,415,000 in funds were distributed, yielding 2,060 responses.
+Survey operations continued until \$2,415,000 in funds were exhausted, yielding 2,060 responses.
 The survey included 100 questions and was conducted in nine different languages: Arabic, Bangla, Cantonese, English, French, Mandarin, Spanish, Tibetan, and Wolof.
-[^4] Vendors who spoke other languages, such as Dari, Farsi, Russian, and Turkish, typically spoke English with a high level of fluency, and those surveys were conducted in English.
-Similarly, vendors who spoke indigenous languages from Latin America, such as Nahuatl and Quechua, typically spoke Spanish with a high level of fluency.
 The survey items solicited a variety of information from vendors, such as logistical information (e.g., vending location, residential location, and frequency of operation), economic information (e.g., items sold, income, and expenses), and demographic information (e.g., age, race, ethnicity, and immigration status).
@@
 Of the 559 merchandise vendors, 505 were not veterans, of which 308 (61%) indicated they had a license to vend.
-The number of respondents is listed by neighborhood in Section 7.1, Table 1.
-(Neighborhoods with few respondents are grouped together.) A map of the number of respondents by ZIP Code Tabulation Area (ZIP Code) is shown in the top left panel of Section 7.2, Figure 1.
+The number of respondents is listed by neighborhood in Table 1.
+(Neighborhoods with few respondents are grouped together.) A map of the number of respondents by ZIP Code Tabulation Area (ZIP Code) is shown in the top left panel of Figure 1.
 Approximately five percent of vendors did not identify their vending location.
@@
 (For this reason, the Respondent and Population columns do not sum exactly to the New York City total.) Veteran merchandise vendors are excluded entirely from Table 1 and Figure 1, although these vendors are reflected in our overall estimate of 23,000 vendors.
-3. Methodology and Data Analysis
-We estimate approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,400 general merchandise vendors.
-We arrive at the estimate by using ratio estimation, which leverages the fact that the number of licenses and permits are limited by local law.
-In Section 3.1, we provide a simple explanation of ratio estimation, which is intended to be accessible to a general readership.
-In Section 3.2, we provide a theoretical justification of our estimates based on the theory of point processes, which, while more technical, is intended to provide a somewhat general argument for ratio estimation with spatial data when a design-based approach cannot be justified.
-Inference is conducted using increasing domain asymptotics, and the details are outlined in Section 7.3.
-In Section 3.3, we apply the methodology to the survey data discussed in Section 2.2.
+3. Methodology
+We estimate the number of street vendors that operate in New York City.
+We use ratio estimation, which leverages the fact that the number of licenses and permits are limited by local law.
+In Section 3.1, we provide a simple explanation of ratio estimation, which is intended to be accessible to a general readership and highlights our main assumption.
+In Section 3.2, we develop a theoretical justification based on the theory of point processes, which, while more technical, allows us to explicitly state the estimands and underlying assumptions and to derive the standard errors.
+Inference is conducted using increasing domain asymptotics, and the details are outlined in Sections 7.1, 7.2, and 7.3.
 3.1 Ratio Estimation
@@
 The purpose is to highlight the main assumption.
-A more rigorous derivation is provided in Section 3.2.
-Consider a fixed region $A$, and let $N_{i}(A)$ denote the number of individuals in region $A$ of type $i$.
-To fix ideas, let $A$ be New York City, let $i=0$ indicate mobile food vendors with a permit, and let $i=1$ indicate mobile food vendors without a permit.
-Also, let $n_{i}(A)$ denote the number of respondents of type $i$.
-We observe $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(A)$ from which we estimate $\theta(A)=\mathbb{E}\left[N_{0}(A)\right]$, the expected number of vendors without a permit.
-The expected number of mobile food vendors with or without a permit is then $\tau(A)=\theta(A)+N_{1}(A)$.
-The ratio estimator of $\theta(A)$ is
-$\hat{\theta}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}$
+A more rigorous derivation is presented in Section 3.2.
+Consider a fixed region $A$, and let $N_{i}(A)$ denote the number of individuals in region $A$ of status $i$.
+To fix ideas, let $A$ be New York City, let $i=0$ indicate mobile food vendors without a permit, and let $i=1$ indicate mobile food vendors with a permit.
+Also, let $n_{i}(A)$ denote the number of respondents of status $i$.
+We observe $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(A)$ from which we would like to determine $N_{0}(A)$, the number of vendors without a permit.
+The total number of mobile food vendors with or without a permit is then $N_{0}(A)+N_{1}(A)$.
+We define the ratio estimator as
+$\hat{\Lambda}_{0}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}$
 It is called the ratio estimator because it depends on the ratio of random variables $n_{0}(A)\,/\,n_{1}(A)$.
-(The total number of permitted vendors, $N_{1}(A)$, is considered fixed.) The estimator is accurate when the response rate, $p$, is approximately the same for each type, i.e.,
+(The total number of permitted vendors, $N_{1}(A)$, is considered fixed.) The estimator is approximately equal to $N_{0}(A)$ when the response rate, $p$, is approximately the same for each status, i.e.,
 $p\approx\frac{n_{0}(A)}{N_{0}(A)}\approx\frac{n_{1}(A)}{N_{1}(A)}$
-in which case $\hat{\theta}(A)$ is the solution for $N_{0}(A)$ in terms of $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(A)$, obtained by cross-multiplication.
+in which case $\hat{\Lambda}_{0}(A)$ is the solution for $N_{0}(A)$ in terms of $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(A)$, obtained by cross-multiplication.
+It follows that $\hat{\tau}(A)=\hat{\Lambda}_{0}(A)+N_{1}(A)$ is approximately equal to the total $N_{0}(A)+N_{1}(A)$.
 The main assumption of ratio estimation is that the survey data are representative in the sense that the response rate does not depend on whether a mobile food vendor has a permit.
-If we further assume that the response rate does not vary by subregion, we can also estimate the total number of mobile food vendors by subregion.
-For example, suppose the response rate in subregion $B\subset A$ is also approximately equal to $p$, e.g.,
+If we further assume that the response rate does not vary by subregion, we can also determine the total number of mobile food vendors by subregion.
+For example, suppose the response rate in subregion $B\subset A$ is also approximately equal to $p$, i.e.,
 $p\approx\frac{n_{1}(A)}{N_{1}(A)}\approx\frac{n_{0}(B)}{N_{0}(B)}$
-Then we can estimate $\theta(B)$ by the subregion estimator
-$\tilde{\theta}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}$
-where $\tilde{\theta}(B)$ is the solution for $N_{0}(B)$ in terms of $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(B)$.
-The subregion estimator is necessary when $N_{1}(B)$ is not observed and $\hat{\theta}(B)$ cannot be calculated directly.
+Then the subregion estimator
+$\tilde{\Lambda}_{0}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}$
+is the solution for $N_{0}(B)$ in terms of $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(B)$.
+The subregion estimator is useful when $N_{1}(B)$ is not observed and $\hat{\Lambda}_{0}(B)$ cannot be calculated directly.
 This is the case with the survey data described in Section 2.2.
-We observe $N_{1}(A)$, the number of permits city-wide (i.e., within region $A$), but not $N_{1}(B)$, the number of permits within subregion $B$.
-(We also observe $n_{0}(B)$ and $n_{1}(B)$.)
-The subregion estimator $\tilde{\theta}(B)$ works by replacing the term $N_{1}(B)$ in the ratio estimator $\hat{\theta}(B)$ with the estimate
-$\tilde{N}_{1}(B)=\dfrac{N_{1}(A)\,n_{1}(B)}{n_{1}(A)}$
-The estimate $\tilde{N}_{1}(B)$ is also used to estimate the total number of mobile food vendors expected in subregion $B$. That is, we estimate $\tau(B)=\theta(B)+N_{1}(B)$ with
-$\tilde{\tau}(B)=\tilde{\theta}(B)+\tilde{N}_{1}(B)=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(A)}$
+We observe $N_{1}(A)$, the number of permits city-wide, but not $N_{1}(B)$, the number of permits within subregion $B$.
+(Recall that the vending locations of the respondents are recorded and thus $n_{0}(B)$ and $n_{1}(B)$ are observed for any $B\subset A$.)
+The subregion estimator $\tilde{\Lambda}_{0}(B)$ works by replacing the term $N_{1}(B)$ in the ratio estimator $\hat{\Lambda}_{0}(B)$ with
+$\dfrac{N_{1}(A)\,n_{1}(B)}{n_{1}(A)}$
+which is justified when
+$p\approx\frac{n_{1}(A)}{N_{1}(A)}\approx\frac{n_{1}(B)}{N_{1}(B)}$
+This suggests the estimator
+$\tilde{\tau}(B)=\tilde{\Lambda}_{0}(B)+\dfrac{N_{1}(A)\,n_{1}(B)}{n_{1}(A)}=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(A)}$
+is approximately equal to the subregion total $N_{0}(B)+N_{1}(B)$.
 3.2 Model
 We provide a point process justification of the ratio estimation procedure described in Section 3.1.
-The purpose is to more closely examine the underlying assumptions and derive the standard errors of the estimates.
-Let $\Pi_{i}$ denote an inhomogeneous spatial Poisson process referencing the location $x\in A\subset\mathbb{R}^{2}$ of each street vendor with status $i$ at the time the survey was conducted.
-As in Section 3.1, we fix ideas by letting $i=0$ indicate mobile food vendors with a permit and $i=1$ indicate mobile food vendors without a permit.
+The purpose is to explicitly state the estimands and underlying assumptions and to derive the standard errors.
+Let $\Pi_{i}$ denote an inhomogeneous Poisson process referencing the location $x\in A\subset\mathbb{R}^{2}$ of each vendor with status $i$ at the time the survey was conducted.
+As in Section 3.1, we fix ideas by letting $i=0$ indicate mobile food vendors without a permit and $i=1$ indicate mobile food vendors with a permit.
 The Poisson process assumption may be justified by the law of rare events.
-New York City can be partitioned into a large number of theoretically vendable locations.
-Whether a vendor is located within a partition has a vanishingly small probability, such that the number of vendors in any area is well approximated by a Poisson distribution.
-This approximation is accurate even if these probabilities are weakly dependent, see Freedman (1974).
+Any sufficiently large area can be partitioned into a large number of theoretically vendable locations.
+Whether a vendor is located within a partition has a vanishingly small probability, such that the number of vendors in that area is well approximated by a Poisson distribution.
+This approximation is accurate even under weak dependence (Freedman 1974).
 Other justifications are possible.
@@
 […]
+The data are the locations of the respondents by status (i.e., the realizations of $\pi_{i}$) and the number of vendors in region $A$ with a permit, $N_{1}(A)$.
+In other words, we assume $n_{0}(B)$ and $n_{1}(B)$ are observed for any $B\subseteq A$, while $N_{0}(B)$ and $N_{1}(B)$ are unobserved with the exception of $N_{1}(A)$.
+Our goal is to estimate the expected number of street vendors in subregion $B\subseteq A$ without a permit,
 […]
+and the expected total with or without a permit,
+$\tau(B)=\Lambda_{0}(B)+\mathbb{E}\left[N_{1}(B)\,|\,N_{1}(A)\right]$
+We condition on $N_{1}(A)$ to reflect the fact that while the number of vendors of either status in any subregion $B\subsetneq A$ is random, the total number of vendors with permits citywide is fixed by law.
 In Section 3.1, we described the main assumption of ratio estimation: that the status and spatial distribution of the respondents is representative.
 We now make this statement precise.
-We assume that $p_{i}(x)$ can be decomposed into a constant that does not depend on $i$ or $x$ plus a spatially varying error term, $p_{i}(x)=p\,+\,\epsilon_{i}(x)$, and the error term is orthogonal to the corresponding mean measure.
+We assume that $p_{i}(x)$ can be decomposed into a constant that does not depend on $i$ or $x$ plus a spatially varying error term, $p_{i}(x)=p\,+\,\epsilon_{i}(x)$, and the error term is orthogonal to the corresponding intensity.
 i.e.,
@@
 […]
-so that by conditioning on $N_{1}(B)$, we arrive at the following probability model for the number of respondents with and without permits
-$n_{0}(B)\sim$ $\ \text{Poisson}\big{(}p\,\theta(B)\big{)}$
-$n_{1}(B)\ |\ N_{1}(B)\sim$ $\ \text{Binomial}\big{(}p,\,N_{1}(B)\big{)}$
-The maximum likelihood estimates for $p$ and $\theta(A)$ are
-$\hat{p}=\dfrac{n_{1}(A)}{N_{1}(A)}\qquad\text{and}\qquad\hat{\theta}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}$
-Using increasing domain asymptotics, the estimator is asymptotically normal with mean $\theta(A)$ and standard error
-$\text{SE}[\hat{\theta}(A)]=\theta(A)\sqrt{\dfrac{1}{p\,\theta(A)}+\dfrac{1-p}{p\,N_{1}(A)}}$
-Substituting $\hat{p}$ and $\hat{\theta}(A)$ for $p$ and $\theta(A)$ yields the following plug-in estimator for the standard error,
-$\hat{\text{SE}}[\hat{\theta}(A)]=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(A)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
-See Section 7.3.2 for details.
-The maximum likelihood estimate for $\theta(B)$ is
-$\tilde{\theta}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}$
-which is asymptotically normal with mean $\theta(B)$ and standard error
-$\text{SE}[\tilde{\theta}(B)]=\theta(B)\sqrt{\frac{1}{p\,\theta(B)}+\frac{1-p}{p\,N_{1}(A)}}$
-The plug-in estimator for the standard error is
-$\tilde{\text{SE}}[\tilde{\theta}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
-To determine the subregion total $\tau(B)=\theta(B)+N_{1}(B)$, we estimate $\theta(B)$ and $N_{1}(B)$ with $\tilde{\theta}(B)$ and $\tilde{N}_{1}(B)=N_{1}(A)\,n_{1}(B)\,/\,n_{1}(A)$, yielding
-$\tilde{\tau}(B)=\tilde{\theta}(B)+\tilde{N}_{1}(B)=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(A)}=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(B)+n_{1}(A\backslash B)}$
-where $A\backslash B$ denotes the compliment of subregion $B$.
+so that by conditioning on $N_{1}(A)$, we arrive at the following probability model for the number of respondents with and without permits
+$n_{0}(B)\sim$ $\ \text{Poisson}\big{(}p\,\Lambda_{0}(B)\big{)}$
+$n_{1}(B)\,|\,N_{1}(A)\sim$ $\ \text{Binomial}\big{(}N_{1}(B),\,p\,q(B)\big{)}$
+where
+$q(B)=\frac{\Lambda_{1}(B)}{\Lambda_{1}(A)}$
+The maximum likelihood estimates for $\Lambda_{0}(A)$ and $p$ are
+$\hat{\Lambda}_{0}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\qquad\text{and}\qquad\hat{p}=\dfrac{n_{1}(A)}{N_{1}(A)}$
+Using increasing domain asymptotics, $\hat{\Lambda}_{0}(A)$ is asymptotically normal with mean $\Lambda_{0}(A)$ and standard error
+$\hat{\text{SE}}[\hat{\Lambda}_{0}(A)]=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(A)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
+The maximum likelihood estimate for the total $\tau(A)$ is $\hat{\tau}(A)=\hat{\Lambda}_{0}(A)+N_{1}(A)$, which has the same standard error as $\hat{\Lambda}_{0}(A)$. See Section 7.1.2 for details.
+The maximum likelihood estimate for $\Lambda_{0}(B)$ is
+$\tilde{\Lambda}_{0}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}$
+which is asymptotically normal with mean $\Lambda_{0}(B)$ and standard error
+$\tilde{\text{SE}}[\tilde{\Lambda}_{0}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
+Finally, the maximum likelihood estimate for the subregion total $\tau(B)$ is
+$\tilde{\tau}(B)=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(A)}$
 The asymptotic standard error is
-$\text{SE}[\tilde{\tau}(B)]=\theta(B)\sqrt{\frac{1}{p\,\theta(B)}+\frac{1-p}{p\,N_{1}(A)}\left(1+\frac{N_{1}(B)\,N_{1}(A\backslash B)}{\theta(B)^{2}}\right)}$
-and the plug-in estimator is
-$\tilde{\text{SE}}[\tilde{\tau}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\left(\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}\right)\left(1+\dfrac{n_{1}(B)\,n_{1}(A\backslash B)}{n_{0}(B)^{2}}\right)}$
-See Section 7.3.3 for details.
-3.3 Data Analysis
-We use the methodology described in Section 3.1 and 3.2 to estimate the number of street vendors in New York City.
+$\tilde{\text{SE}}[\tilde{\tau}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\frac{1}{n_{1}(A)}-\frac{1}{N_{1}(A)}+\dfrac{n_{1}(B)\,n_{1}(A\backslash B)}{n_{1}(A)\,n_{0}(B)^{2}}}$
+where $A\backslash B$ denotes the complement of subregion $B$. See Section 7.1.3 for details.
+4. Model Validation
+We make two assumptions in Section 3.
+The main assumption is that the status and spatial distribution of the respondents is representative.
+This assumption is necessary to ensure the ratio estimator is consistent.
+A secondary assumption is that the spatial distribution of vendors is well described by a family of inhomogeneous Poisson processes.
+This assumption is necessary to derive the standard errors.
+In this section, we relax both assumptions.
+Since neither can be relaxed from the data alone, we consider additional information.
+In Section 4.1, we examine a generalization of the ratio estimator in which sampling weights adjust for differential nonresponse.
+In Section 4.2, we model the arrival of street vendors as a pure birth process, and we derive a more general formula for the standard errors in which the amount of extra-Poisson variation can be determined by the number of vendors that cluster within markets.
+4.1 Representativeness
+The main assumption in Section 3 is that the status and spatial distribution of the respondents is representative.
+If this assumption holds, then the ratio estimator is consistent for a wider class of point processes in which Campbell’s Theorem applies, not just the Poisson process.
+See Daley and Vere-Jones (2007) for a general statement of Campbell’s Theorem.
+If this assumption is violated, then $\hat{\Lambda}_{0}(B)$ may be inconsistent.
+For example, if respondents are more likely to come from locations in which a higher proportion of vendors have permits, then the ratio $n_{0}(B)\,/\,n_{1}(B)$ would not be representative, and $\hat{\Lambda}_{0}(B)$ would likely underestimate $\Lambda_{0}(B)$, even in large samples.
+We think the representativeness assumption is reasonable because the Street Vendor Project canvassed New York City to distribute a large amount of relief aid as discussed in Section 2.
+Nevertheless, to evaluate how the estimates might change when this assumption is violated, we generalize the estimator in Section 3.2 to accommodate spatially varying sampling weights.
+The generalization is useful for determining the sensitivity of our estimates to differential nonresponse as we demonstrate in Section 5.1.
+Let $w_{0}(x)$ and $w_{1}(x)$ denote two spatially varying weight functions. We define the weighted counts
+$n^{w}_{0}(B)=\sum_{x\,\in\,\pi_{0}\cap B}w_{0}(x)\qquad\text{and}\qquad n^{w}_{1}(B)=\sum_{x\,\in\,\pi_{1}\cap B}w_{1}(x)$
+where $\pi_{i}\cap B$ refers to the points of the process $\pi_{i}$ that fall in subregion $B\subseteq A$. We also define the weighted estimators for $\Lambda_{0}(A)$ and $p$,
+$\hat{\Lambda}_{0}^{w}(A)=\dfrac{N_{1}(A)\,n^{w}_{0}(A)}{n^{w}_{1}(A)}\qquad\text{and}\qquad\hat{p}^{w}=\frac{n^{w}_{1}(A)}{N_{1}(A)}$
+Under Campbell’s Theorem, the means $\mathbb{E}[n^{w}_{0}(A)]$ and $\mathbb{E}[n^{w}_{1}(A)|N_{1}(A)]$ are
+[…]
+We assume the weight functions $w_{0}$ and $w_{1}$ can be chosen in such a way that the expected value is inversely proportional with the probability of a response. That is,
+$p=\mathbb{E}[w_{0}(x)]\,p_{0}(x)=\mathbb{E}[w_{1}(x)]\,p_{1}(x)$
+It follows that $\mathbb{E}[n^{w}_{0}(A)]$ and $\mathbb{E}[n^{w}_{1}(A)\,|\,N_{1}(A)]$ simplify to
+$\mathbb{E}\left[n^{w}_{0}(A)\right]=p\,\Lambda_{0}(A)\qquad\text{and}\qquad\mathbb{E}\left[n^{w}_{1}(A_{i})\,|\,N_{1}(A)\right]=p\,N_{1}(A)$
+and that $\hat{\Lambda}_{0}^{w}(A)$ is approximately normal with mean $\Lambda_{0}(A)$ and standard error
+$\hat{\text{SE}}[\hat{\Lambda}_{0}^{w}(A)]=\dfrac{N_{1}(A)\,n_{0}^{w}(A)}{n_{1}^{w}(A)}\sqrt{\dfrac{\hat{\mu}^{2}_{0}(A)+\hat{\sigma}_{00}(A,A)}{n_{0}^{w}(A)^{2}}+\dfrac{\hat{\mu}_{1}^{2}(A)}{n^{w}_{1}(A)^{2}}+\dfrac{(N_{1}(A)-1)\,\hat{\sigma}_{11}(A,A)}{N_{1}(A)\,n^{w}_{1}(A)^{2}}-\dfrac{1}{N_{1}(A)}-\dfrac{2\,\hat{\sigma}_{01}(A,A)}{n_{0}^{w}(A)\,n_{1}^{w}(A)}}$
+where
+$\hat{\mu}_{k}^{2}(A)=\sum_{x\,\in\,\pi_{k}\cap A}\mathbb{E}\bigl{[}w_{k}^{2}(x)\bigr{]}\qquad\text{and}\qquad\hat{\sigma}_{kl}(A,A)=\sum_{\begin{subarray}{c}x\,\in\,\pi_{k}\cap A\\[1.39998pt] y\,\in\,\pi_{l}\cap A\\[1.39998pt] x\neq y\end{subarray}}\mathbb{C}\bigl{(}w_{k}(x),\,w_{l}(y)\bigr{)}$
+See Section 7.1.4 for details including derivations for the subregion and subtotal estimators.
+The standard error derived in this section for $\hat{\Lambda}_{0}(A)$ is identical to the standard error $\hat{\Lambda}_{0}(A)$ from Section 3.2 if $w_{0}(x)=w_{1}(x)=1$.
+In this case, $\hat{\mu}_{0}^{2}(A)=n_{0}(A)=n_{0}^{w}(A)$, $\hat{\mu}_{1}^{2}(A)=n_{1}(A)=n^{w}_{1}(A)$, and $\hat{\sigma}_{00}=\hat{\sigma}_{01}=\hat{\sigma}_{11}=0$.
+Note that $\hat{\Lambda}_{0}(A)$ and $\hat{\Lambda}_{0}^{w}(A)$ are related by the relationship
+$\hat{\Lambda}_{0}^{w}(A)=\frac{N_{1}(A)\ n^{w}_{0}(A)}{n^{w}_{1}(A)}=\hat{\Lambda}_{0}(A)\,\frac{n^{w}_{0}(A)\,/\,n_{0}(A)}{n^{w}_{1}(A)\,/\,n_{1}(A)}$
+It follows that the relative bias depends only on the ratio of the average weights $n^{w}_{0}(A)\,/\,n_{0}(A)$ and $n^{w}_{1}(A)\,/\,n_{1}(A)$.
+For example, if the average permitted vendor is twice as likely to respond as the average unpermitted vendors, $n^{w}_{0}(A)\,/\,n_{0}(A)$ is twice the size of $n^{w}_{1}(A)\,/\,n_{1}(A)$, and $\hat{\Lambda}_{0}(A)$ is twice the size of $\hat{\Lambda}_{0}^{w}(A)$.
+The estimators $\hat{\tau}(A)$ and $\hat{\tau}^{w}(A)$ are similarly related by the relationship
+$\hat{\tau}^{w}(A)=\frac{N_{1}(A)\left(n^{w}_{0}(A)+n^{w}_{1}(A)\right)}{n^{w}_{1}(A)}=\hat{\tau}(A)\,\frac{\left(n^{w}_{0}(A)+n^{w}_{1}(A)\right)\,/\,\left(n_{0}(A)+n_{1}(A)\right)}{n^{w}_{1}(A)\,/\,n_{1}(A)}$
+The same is true for subregion and subtotal estimators, see Section 7.1.4.
+4.2 Overdispersion
+The second assumption in Section 3.2 is that the spatial distribution of vendors is well approximated by a family of inhomogeneous Poisson processes.
+We consider this assumption secondary because $\hat{\Lambda}_{0}(A)$ is consistent under the wide class of point processes described in Section 4.1.
+That said, the standard errors derived in Section 3.2 do not necessarily hold when the Poisson assumption is violated.
+We think the Poisson assumption is reasonable because a wide class of data generating processes are well-approximated by the Poisson process under the law of rare events as discussed in Section 3.2.
+Nevertheless, to evaluate how the standard errors might change when this assumption is violated, we consider a general model that describes how street vendors arrived at their present locations.
+Specifically, we consider two families of pure birth processes: one with a constant arrival rate (or “birth rate”) in which the Poisson assumption holds and one with a linear arrival rate in which the Poisson assumption does not hold.
+Suppose vendors serve customers who congregate at fixed locations, which we call markets.
+Let $M(B)$ denote the number of markets in any subregion $B\subseteq A\subset\mathbb{R}^{2}$.
+If every market is in a competitive equilibrium such that the expected profit for a new vendor is the same at each market, then the time between vendor arrivals at each market may be well-approximated by an exponential distribution with a constant arrival rate $\theta_{i}$.
+Further suppose each market starts with one vendor at $t=0$.
+Then at the time the survey was conducted ($t=1$), the number of vendors, $N_{i}(B)$, would follow a Poisson distribution with mean $\Lambda_{i}(B)=\theta_{i}\,M(B)$, and the results of Section 3.2 hold.
+Now suppose instead that larger markets grow faster than smaller markets, an empirical phenomenon known as preferential attachment.
+Then the time between vendor arrivals may be better approximated by an exponential distribution with a linear arrival rate, or Yule process.
+That is, let $N^{t}_{i}(B)$ denote the number of vendors in subregion $B$ at time $t$ so that the time between vendor arrivals at each market follows an exponential distribution with rate $\theta_{i}\,N^{t}_{i}(B)$.
+If we assume further that each market started with one vendor at $t=0$, then the number of vendors at the time the survey was conducted ($t=1$), $N_{i}(B)$, would follow a negative binomial distribution with mean $\Lambda_{i}(B)=\exp(\theta_{i})\,M(B)$ and variance $\Lambda_{i}(B)\bigr{(}\Lambda_{i}(B)-M(B)\bigr{)}\,/\,M(B)$.
+i.e.,
+$N_{i}(B)\sim\ \text{Negative Binomial}\left(\Lambda_{i}(B),\ \Lambda_{i}(B)\,\frac{\Lambda_{i}(B)-M(B)}{M(B)}\right)$
+where we parameterize the negative binomial distribution by its mean and variance for comparison with Section 3.2.
+In this case, $N_{i}(B)$ exhibits extra-Poisson variation, and the results of Section 3.2 do not hold.
+See Ross (2014, chap.
+6) for details.
+To derive the standard errors, suppose a portion of the markets, $p$, were selected for the survey at random so that
+$n_{i}(B)\sim\ \text{Negative Binomial}\left(p\,\Lambda_{i}(B),\ p\,\Lambda_{i}(B)\,\frac{\Lambda_{i}(B)-M(B)}{M(B)}\right)$
+The conditional distribution of $n_{1}(B)$ given $N_{1}(A)$ is then
+$n_{1}(B)\,|\,N_{1}(A)\sim\ \text{Negative Hypergeometric}\bigl{(}N_{1}(A)-1,\ M(A)-1,\ p\,q(B)\,M(A)\bigl{)}$
+The conditional mean and variance of $n_{1}(A)$ given $N_{1}(A)$ is
+$\mathbb{E}\left[n_{1}(A)\,|\,N_{1}(A)\right]=p\,N_{1}(A)\qquad\text{and}\qquad\mathbb{V}\bigr{(}n_{1}(A)\,|\,N_{1}(A)\bigr{)}=u_{1}(A)\,p\,(1-p)\,N_{1}(A)$
+so that under increasing domain asymptotics, the ratio estimator
+$\hat{\Lambda}_{0}(A)=\frac{N_{1}(A)\ n_{0}(A)}{n_{1}(A)}$
+is asymptotically normal with mean $\Lambda_{0}(A)$ and standard error
+$\hat{\text{SE}}[\hat{\Lambda}_{0}(A)]=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\sqrt{\hat{u}_{0}(A)\left(\dfrac{1}{n_{0}(A)}\right)+u_{1}(A)\left(\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}\right)}$
+where
+$\hat{u}_{0}(B)=\cfrac{\dfrac{N_{1}(A)\ n_{0}(B)}{n_{1}(A)}-M(B)}{M(B)}\qquad\text{and}\qquad u_{1}(A)=\dfrac{N_{1}(A)-M(A)}{M(A)+1}$
+See Section 7.1.5 for details including derivations for the subregion and subtotal estimators. Note that our parameterization makes use of the relationship $q(B)=M(B)\,/\,M(A)$.
+The standard error derived in this section for $\hat{\Lambda}_{0}(A)$ is identical to the standard error for $\hat{\Lambda}_{0}(A)$ from Section 3.2 except for the weights $u_{0}$ and $u_{1}$.
+These weights are the result of the extra-Poisson variation induced by preferential attachment—that is, the linear “birth” rate in which the arrival of new vendors is proportional to the number of vendors in the market.
+For example, if there are five vendors in each market on average, then $u_{0}$ and $u_{1}$ are approximately 4 and the standard error of the ratio estimator derived in this section is twice the size of the standard error derived in Section 3.2.
+The same is true for subregion and subtotal estimators, see Section 7.1.5.
+5. Data Analysis
+We apply the results presented in Section 3 and 4 to the data described in Section 2.
+In Section 5.1, we use the ratio estimator described in Section 3.1 and 3.2 to estimate the number of street vendors in New York City.
+In Section 5.2, we use the extensions described in Section 4.1 and 4.2 to deteremine the sensitivity of the estimate to the underlying assumptions.
+Sensitivity is assessed using additional data from the NYC Office of Administrative Trials and Hearings (OATH) (accessed 2024-04-03) and the American Community Survey from the U.S.
+Census Bureau (accessed 2022-12-22).
+5.1 Estimation
 According to the survey data described in Section 2.2, $n_{1}(A)=349$ mobile food vendor respondents indicated they had a permit, and $n_{0}(A)=1,051$ indicated they did not.
@@
 When combined with the estimated $1,000$ veteran merchandise vendors (see Section 2.1) and $20,500$ food vendors, the total number of street vendors is approximately $23,000$.
-Estimates for select subregions are listed in Section 7.1, Table 1.
+Estimates for select subregions are listed in Table 1.
 We find that approximately one quarter of street vendors are located in the Manhattan neighborhoods of Chelsea, Clinton, and Lower Manhattan.
@@
 We also report the margin of error (i.e., two standard errors, half of the width of a 95% confidence interval).
-We check these estimates by comparing them to a second, independent assessment conducted by the Street Vendor Project in the Bronx on May 13 and 15, 2022, in which the number of street vendors was documented through on-the-ground observations.
+These estimates can be checked by comparing them to a second, independent assessment conducted by the Street Vendor Project in the Bronx on May 13 and 15, 2022, in which the number of street vendors was documented through on-the-ground observations.
 The independent assessment identified 188 total street vendors near Fordham Road (Fordham Road BID and Street Vendor Project 2024).
-According to the survey data, 17 respondents indicated that they operate near Fordham Road, yielding an estimate of 152 street vendors total with a standard error of 54.
+According to the survey data, 17 respondents indicated that they operate near Fordham Road, yielding an estimate of 166 street vendors total with a standard error of 54.
 We conclude this estimate is consistent with the 188 street vendors counted independently by Fordham Road BID and Street Vendor Project (2024).
-4. Model Validation
+{ThreePartTable}{TableNotes}
+This table provides estimates for the number of mobile food vendors and non-veteran general merchandise vendors.
+The number of veteran general merchandise vendors is approximately 1,000.
+First amendment vendors and vendors in NYC Parks are not included.
+Table 1: Number of Street Vendors
+Respondents Population Margin of Error
+Bronx
+   Bronx Park and Fordham 68 757 239
+   Southeast Bronx 28 385 160
+   High Bridge and Morrisania 20 233 127
+   Central Bronx 9 96 81
+   Other Bronx 31 264 137
+   Total 156 1,735 376
+Brooklyn
+      Northwest Brooklyn 46 613 204
+      Bushwick and Williamsburg 41 433 175
+      Sunset Park 26 344 153
+                  Flatbush 19 242 128
+            Borough Park 10 134 94
+            Southwest Brooklyn 6 64 66
+            Other Brooklyn 28 374 159
+            Total 176 2,205 418
+Manhattan
+            Chelsea and Clinton 341 3,289 490
+         Lower Manhattan 105 1,250 287
+         Gramercy Park and Murray Hill 97 1,062 274
+         Lower East Side 65 808 232
+         Greenwich Village and Soho 61 726 222
+         Upper East Side 56 534 194
+         Central Harlem 46 483 185
+         Inwood and Washington Heights 33 459 176
+         Upper West Side 48 429 172
+         East Harlem 38 425 174
+         Total 890 9,464 879
+Queens
+         West Queens 396 4,650 692
+      North Queens 67 896 251
+      Northwest Queens 23 312 144
+      Jamaica 14 181 110
+      Other Queens 31 406 168
+      Total 531 6,445 860
+Staten Island
+      Total 4 58 58
+New York City
+   Total 1,905 21,857 1,941
+[…]
+Figure 1: Location of Street Vendors
+[Uncaptioned image][Uncaptioned image]
+[Uncaptioned image][Uncaptioned image]
+These maps compare the spatial distribution of respondents (left) to the distribution of vendors suggested by administrative data (New York City Office of Administrative Trials and Hearings (OATH), top right) and federal data (American Community Survey (ACS), bottom right).
+5.2 Model Validation
 We make two assumptions in Section 3.
 The main assumption is that the status and spatial distribution of the respondents is representative.
-This assumption is necessary to ensure the ratio estimator is accurate in large samples (i.e., consistent).
-A secondary assumption is that the spatial distribution of vendors is well described by a family of inhomogeneous spatial Poisson processes.
-This assumption is necessary to derive the standard errors.
-In this section, we examine each assumption.
-Since neither can be evaluated from the survey data alone, we consider additional information.
-In Section 4.1, we compare the spatial distribution of the respondents to data from the New York City Office of Administrative Trials and Hearings and the American Community Survey.
-In Section 4.2, we model the arrival of street vendors as a pure birth process, and we derive a more general formula for the standard errors in which the amount of extra-Poisson variation can be determined by the number of vendors that cluster within markets.
-[^5] The results of this section can also be derived from a spatial Negative Binomial Lévy process in a manner that mirrors Section 3.2.
-The advantage of the birth process approach taken in Section 4.2 is that it provides a plausible explanation for why the Poisson assumption made in Section 3.2 may not hold in practice.
-4.1 Representativeness
-The main assumption in Section 3 is that the status and spatial distribution of the respondents is representative (i.e., the response probability is spatially uncorrelated with the mean measure).
-If this assumption holds, then the ratio estimator is consistent for a wider class of point processes in which Campbell’s Theorem applies, not just the Poisson process.
-See Daley and Vere-Jones (2007) for a general statement of Campbell’s Theorem.
-If this assumption is violated, then $\hat{\theta}(A)$ may be inconsistent.
-For example, if respondents are more likely to come from locations in which a higher proportion of vendors have permits, then the ratio $n_{0}(A)\,/\,n_{1}(A)$ would not be representative, and $\hat{\theta}(A)$ would likely underestimate $\theta(A)$ even in large samples.
-We think the representativeness assumption is reasonable because the Street Vendor Project canvassed New York City to distribute a large amount of relief aid.
-We argue that trust in SVP and its mission, along with the aid, removed many barriers that typically prevent survey operations from reaching hard-to-reach populations.
+We believe trust in SVP and its mission, along with the aid, removed many barriers that typically prevent survey operations from reaching hard-to-reach populations.
 Instead, nonresponse reflects chance variation in the personal circumstances of a vendor that are largely unrelated to location.
-Nevertheless, it is possible that canvassing operations systematically missed locations, and these missed locations may be revealed by examining other data sources.
+Nevertheless, it is possible that canvassing operations systematically missed locations.
 We thus compare the spatial distribution of vending locations reported by respondents to the spatial distribution of vendors who allegedly violated street vending laws during the year 2021.
@@
 Census Bureau (accessed 2022-12-22).
-These comparisons are visualized in Section 7.2, Figure 1.
-We find the spatial distribution of vending locations largely agree (top panels) as do the spatial distribution of the residencies (bottom panels), although there are several small discrepancies.
-For example, the OATH data suggest a larger percentage of vendors work in South Brooklyn (Coney Island and Brighton Beach) than captured by the Street Vendor Project survey, and the ACS data suggest a larger percentage of street vendors live in South Queens (Rockaway).
-However, both the OATH and ACS data suggest these locations reflect only a small percentage of New York City vendors, and thus any resulting undercount is likely to have little impact on our overall estimate.
-We also note that even if this discrepancy suggests a location was missed by the Street Vendor Project, our main assumption is not necessarily violated.
-That is because the undercount in some areas may be compensated by an overcount in others (e.g., the Upper East Side of Manhattan or East Brooklyn) such that our overall estimate is still accurate.
-4.2 Overdispersion
-The second assumption in Section 3.2 is that the spatial distribution of vendors is well approximated by a family of inhomogeneous Poisson processes.
-We consider this assumption secondary because the consistency of $\hat{\theta}(A)$ holds under a wide class of point processes described in Section 4.1.
-That said, the standard errors derived in Section 3.2 do not necessarily hold when the Poisson assumption is violated.
-[^6] The ratio estimator may no longer be asymptotically efficient as well.
-In Section 7.3, we show that under the Poisson process assumption, the ratio estimator maximizes the likelihood, which implies it is asymptotically efficient.
-However, the ratio estimator may not maximize the likelihood when the Poisson process assumption is violated.
-We think the Poisson assumption is reasonable because a wide class of data generating processes are well-approximated by the Poisson process under the law of rare events as discussed in Section 3.2.
-Nevertheless, to evaluate how the standard errors might change when this assumption is violated, we consider a general model that describes how street vendors arrived at their present locations.
-Specifically, we consider two families of pure birth processes: one with a constant arrival rate (or “birth rate”) in which the Poisson assumption holds and one with a linear arrival rate in which the Poisson assumption does not hold.
-[…]
-Now suppose instead that larger markets grow faster than smaller markets, an empirical phenomenon known as preferential attachment.
-Then the time between vendor arrivals may be better approximated by an exponential distribution with a linear arrival rate, or Yule process (Yule 1925).
-That is, let $N^{t}_{i}(B)$ denote the number of vendors in subregion $B$ at time $t$ so that the time between vendor arrivals at each market follows an exponential distribution with rate $\lambda_{i}\,N^{t}_{i}(B)$.
-If we assume further that each market started with one vendor at $t=0$, then the number of vendors at the time the survey was conducted ($t=1$), $N_{i}(B)$, would follow a negative binomial distribution with mean $\mu_{i}(B)=\exp(\lambda_{i})\,M(B)$ and variance $\mu_{i}(B)\bigr{(}\mu_{i}(B)-M(B)\bigr{)}\,/\,M(B)$.
-i.e.,
-$N_{i}(B)\sim\ \text{Negative Binomial}\left(\mu_{i}(B),\ \mu_{i}(B)\,\frac{\mu_{i}(B)-M(B)}{M(B)}\right)$
-where we parameterize the negative binomial distribution by its mean and variance.
-In this case, $N_{i}(B)$ exhibits extra-Poisson variation, and the results of Section 3.2 do not hold.
-See Ross (2014, chap.
-6) for details.
-[^7] Note that we define the negative binomial and negative hypergeometric distributions as the total number of trials required to get a predetermined number of successes.
-The negative binomial assumes sampling with replacement, while the negative hypergeometric distribution assumes sampling without replacement.
-To derive the standard errors, suppose a portion of the markets, $p$, were selected for the survey so that
-$n_{i}(B)\sim\ \text{Negative Binomial}\left(p\,\mu_{i}(B),\ p\,\mu_{i}(B)\,\frac{\mu_{i}(B)-M(B)}{M(B)}\right)$
-The conditional distribution of $n_{1}(B)$ given $N_{1}(B)$ is then
-$n_{1}(B)\ |\ N_{1}(B)\sim\ \text{Negative Hypergeometric}\bigl{(}N_{1}(B)-1,\ M(B)-1,\ p\,M(B)\bigl{)}$
-with conditional mean and variance
-$\mathbb{E}\left[n_{1}(B)\ |\ N_{1}(B)\right]=p\,N_{1}(B)\qquad\text{and}\qquad\mathbb{V}\bigr{(}n_{1}(B)\ |\ N_{1}(B)\bigr{)}=w_{1}\,p\,(1-p)\,N_{1}(B)$
-so that under increasing domain asymptotics, the ratio estimator
-$\hat{\theta}(A)=\frac{N_{1}(A)\ n_{0}(A)}{n_{1}(A)}$
-is asymptotically normal with mean $\theta(A)=\mu_{0}(A)$ and standard error
-$\text{SE}[\hat{\theta}(A)]=\theta(A)\sqrt{w_{0}\left(\frac{1}{p\,\theta(A)}\right)+w_{1}\left(\frac{1-p}{p\,N_{1}(A)}\right)}$
-where
-$w_{0}=\dfrac{\theta(A)-M(A)}{M(A)}\qquad\text{and}\qquad w_{1}=\dfrac{N_{1}(A)-M(A)}{M(A)+1}$
-See Section 7.3.4 for details. Note that the weights $w_{0}$ and $w_{1}$ depend on $A$, although this dependence is not reflected in the notation.
-The standard error derived in this section for $\hat{\theta}(A)$ is identical to the standard error for $\hat{\theta}(A)$ from Section 3.2 except for the weights $w_{0}$ and $w_{1}$.
-These weights are the result of the extra-Poisson variation induced by preferential attachment—that is, the linear “birth” rate in which the arrival of new vendors is proportional to the number of vendors in the market.
-The weights $w_{0}$ and $w_{1}$ correspond to the relative number of vendors without permits and with permits per market respectively.
+These comparisons are visualized in Figure 1.
+We find the spatial distribution of vending locations largely agree (top panels) as do the spatial distribution of the residencies (bottom panels).
+This suggests that, if there were a region of street vendors missed by the SVP, it was also missed by law enforcement and the ACS.
+We can also combine the data from OATH and ACS with the results from Section 4.1 to assess the sensitivity of the estimates in Section 5.1 to differential nonresponse, as we now demonstrate.
+While the difference between these new estimates and the estimates provided in Section 5.1 are not statistically significant at the 95% level, we believe they provide insight into how the estimates in Section 5.1 might theoretically err.
+First, we consider the scenario in which the probability a vendor responds is inversely proportional to the average number of vendors cited for violating street vending laws in the surrounding ZIP Code each day, as reported in the OATH data for the year 2021.
+This might be the case if vendors in high enforcement areas are reluctant to interact with survey takers.
+The weights are the inverse probabilities—the average number of enforcement actions taken against vendors in the surrounding ZIP Code.
+We find the weighted counts for food vendors are $n^{w}_{0}(A)=25.7$ and $n^{w}_{1}(A)=5.24$.
+This suggests there are $\left(n_{0}^{w}(A)+n_{1}^{w}(A)\right)\,/\,\left(n_{0}(A)+n_{1}(A)\right)\,/\left(n_{1}^{w}(A)\,/\,n_{1}(A)\right)=47\%$ more food vendors than estimated in Section 5.1.
+A similar argument for merchandise vendors suggests there are 9% more merchandise vendors than estimated in Section 5.1.
+Put together, the weights suggest the number of street vendors in New York City is 40% higher than estimated in Section 5.2.
+The second scenario we consider is one in which the probability a vendor responds in proportional to the average number of vendors cited for violating street vending laws in the surrounding ZIP Code.
+This might be the case if vendors in high enforcement areas are easier to identify and thus easier to survey.
+Under this scenario, there are 8% fewer food vendors than estimated in Section 5.1 and 9% fewer merchandise vendors, suggesting the number of street vendors in New York City could be 8% lower than estimated in Section 5.2.
+The final two scenarios are identical to the first two scenarios, except we use the number of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers in the surrounding Public Use Microdata Area as estimated from the 2017-2021 U.S.
+Census Bureau (accessed 2022-12-22).
+The result is less extreme than using the average number of vendors cited.
+If the probability of response is inversely proportional to this estimate, there are 33% more vendors New York City than estimated in Section 5.2.
+If the probability of response is proportional to this estimate, there are 5% fewer vendors.
+Combining these four scenarios suggests a conservative range of between 21,000 and 32,200 street vendors operate in New York City.
+The second assumption is that the spatial distribution of vendors is well described by a family of inhomogeneous spatial Poisson processes.
+We argue that this assumption is reasonable because a wide class of data generating processes are well-approximated by the Poisson process under the law of rare events as discussed in Section 3.2.
+It is possible, however, that vendors cluster within markets as described in Section 4.2.
+The weights $u_{0}$ and $u_{1}$ in Section 4.2 correspond to the relative number of vendors without permits and with permits per market respectively.
 These weights cannot be estimated from the survey data alone.
-However, we believe that $\theta(A)$ and $N_{1}(A)$ are no more than 10 times larger than $M(A)$.
-It follows that $w_{0}$ and $w_{1}$ are less than 9, and thus the standard errors from Section 3.2 and the margins of error from Table 1 are at most 3 times larger if the Poisson assumption is violated and the Yule process holds.
+However, we believe that $\Lambda_{0}(A)$ and $N_{1}(A)$ are no more than 10 times larger than $M(A)$.
+It follows that $u_{0}$ and $u_{1}$ are less than 9, and thus the standard errors from Section 3.2 and the margins of error from Table 1 are at most 3 times larger if the Poisson assumption is violated and the Yule process holds.
 We suspect this bound is conservative, however, and the typical market contains only one or two vendors.
-The same behavior is exhibited by the subregion estimator of $\theta(B)$,
-$\tilde{\theta}(B)=\dfrac{N_{1}(A)\ n_{0}(B)}{n_{1}(A)}$
-which is asymptotically normal with mean $\theta(B)=\mu_{0}(B)$ and standard error
-$\text{SE}[\tilde{\theta}(B)]=\theta(B)\sqrt{v_{0}\left(\frac{1}{p\,\theta(B)}\right)+w_{1}\left(\frac{1-p}{p\,N_{1}(A)}\right)}$
-where
-$v_{0}=\dfrac{\theta(B)-M(B)}{M(B)}$
-The estimator for the subregion total, $\tilde{\tau}(B)=\tilde{\theta}(B)+\tilde{N}_{1}(B)$, is also asymptotically normal with mean $\tau(B)$ and standard error
-$\text{SE}[\tilde{\tau}(B)]=\theta(B)\sqrt{v_{0}\left(\frac{1}{p\,\theta(B)}\right)+\left(\frac{v_{1}\,N_{1}(B)\left(N_{1}(A\backslash B)-\theta(B)\right)^{2}+v_{2}\,N_{1}(A\backslash B)\left(N_{1}(B)+\theta(B)\right)^{2}}{N_{1}(A)\,\theta(B)^{2}}\right)\left(\frac{1-p}{p\,N_{1}(A)}\right)}$
-where
-$v_{1}=\dfrac{N_{1}(B)-M(B)}{M(B)+1}\qquad\text{and}\qquad v_{2}=\dfrac{N_{1}(A\backslash B)-M(A\backslash B)}{M(A\backslash B)+1}$
-See Section 7.3.4 for details.
 6. Discussion
-We conclude by discussing the implications of our estimates.
+We discuss the implications of our estimates.
 In particular, we assess the completeness of the New York City Office of Administrative Trials and Hearings (OATH) and American Community Survey (ACS) data.
@@
 Using the lower limit of a 95% confidence interval for our estimate (21,000) and the upper 95% confidence limit for the ACS estimate (5,610), our findings suggest the ACS misses roughly three quarters of street vendors or more.
-Note the importance of the assumptions stated in Section 3.2 since other derivations of ratio estimation may imply smaller or larger standard errors, from which we might conclude the ACS misses a smaller or larger portion of the street vending population.
-For example, if the standard errors from Section 4.2 hold, then the confidence intervals would overlap if there were more than 300 vendors in the average market.
-We believe this is highly unlikely, suggesting the two estimates are inconsistent even under overdispersion.
+Note the importance of the assumptions stated in Section 3.2 since deviations might suggest the ACS misses a smaller or larger portion of the street vending population.
+For example, if the average unpermitted or unlicensed vendor is 10 times more likely to respond than the average permitted or licensed vendor, then the ACS estimate may miss less than half of the street vending population.
+For another example, if the standard errors from Section 4.2 hold, then the confidence intervals would overlap if there were more than 300 vendors in the average market.
+We believe it is highly unlikely that our assumptions are violated to this extreme, suggesting the two estimates are inconsistent even under differential nonresponse and overdispersion.
 Additional evidence comes from the OATH data, which indicate approximately 2,500 (unique) vendors are summoned to court each year for violating vending laws.
@@
 The fact that the Street Vendor Project distributed relief aid may have encouraged respondents who are unlikely to respond to the ACS.
-Moreover, the Street Vendor Project, through membership lists, referrals, and canvassing may have covered individuals missed by the Census Bureau’s sampling frame.
+Moreover, the Street Vendor Project, through membership lists, referrals, and canvassing may have covered individuals whose households were missed by the Census Bureau’s sampling frame.
+Indeed, a street vendor must be sufficiently visible while on the street in order to sell their items, while the residency of that same vendor may be more easily overlooked.
 Another explanation is that the Street Vendor Project survey and the ACS may be measuring different populations or concepts.
@@
 It is possible that the Street Vendor Project’s definition of a vendor is more inclusive.
-Street vending often provides supplemental income, and thus the discrepancy may reflect the broader challenge of studying the gig economy.
-In this case, either estimate may be relevant, depending on the intended use case.
+Street vending often provides supplemental income, and thus the discrepancy may reflect the broader challenge of studying the “gig economy.” In this case, either estimate may be relevant, depending on the intended use case.
+We believe this application highlights the value of modeling respondents as a point process.
+We believe our approach is relevant for a wide variety of surveys in which units cannot be chosen at random from a sampling frame.
+For example, individuals may be sampled on a beach or at various entrances of a park.
+Future work might explore additional point process models that describe the behavior of the population.
+For example, we have found that when vendors cluster within markets, standard errors become overdispersed.
+However, it is possible that vendors protect their turf, leading to underdispersion.
+It may even be that some vendors benefit from operating near each other, while others benefit from maintaining their distance.

### v6

@@
-Shamier Settle, Matthew Shapiro, Emerita Torres, and Anand Vidyashankar for discussion and feedback on the work described
+Estimating the Number of Street Vendors in New York City:
 Abstract
 We estimate the number of street vendors in New York City.
-We first summarize the process by which vendors receive licenses and permits to legally operate in New York City.
+First, we summarize the process by which vendors receive licenses and permits to operate legally in New York City.
 We then describe a survey that was administered by the Street Vendor Project while distributing coronavirus relief aid to vendors operating in New York City both with and without a license or permit.
-Finally, we calculate the total number of vendors using ratio estimation.
-We find that approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,300 general merchandise vendors.
-One third are located in just six ZIP Codes: 11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan.
-We also provide a theoretical justification of our estimates based on the theory of point processes and a discussion of their accuracy and implications.
-In particular, our estimates suggest the American Community Survey fails to cover the majority of New York City street vendors.
+Finally, we review ratio estimation and develop a theoretical justification based on the theory of point processes.
+We find approximately 23,000 street vendors operate in New York City—20,500 mobile food vendors and 2,400 general merchandise vendors—with one third located in just six ZIP Codes—11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan.
+Our estimates suggest the American Community Survey misses the majority of New York City street vendors.
 1. Introduction
 Street vendors are New York City’s smallest businesses, selling food and merchandise from carts, stalls, and trucks throughout the five boroughs.
-They are an iconic part of the urban landscape and a thriving sector of the local economy, contributing millions of dollars in government revenue through taxes, fines, and fees.
+They are an iconic part of the urban landscape.
+They are also a thriving sector of the local economy, contributing millions of dollars in government revenue through taxes, fines, and fees.
 Perhaps most importantly, street vending historically benefits underserved communities, both because vendors operate in neighborhoods with limited access to traditional stores and because vending is one of a handful of occupations in which New Yorkers of all backgrounds, immigrants in particular, can achieve economic mobility and a chance at the American dream (Burrows and Wallace 1998, chap.
@@
 In Section 3, we use the fact that the number of licenses and permits are limited by law to construct a ratio estimator for the number of street vendors.
-We then develop a theoretical justification based on the theory of point processes, which assumes the spatial distribution of survey respondents is representative and well-approximated by a set of inhomogeneous Poisson processes.
+We then develop a theoretical justification based on the theory of point processes, which assumes the spatial distribution of survey respondents is representative and well approximated by a set of inhomogeneous Poisson processes.
 Finally, we derive the standard errors.
 In Section 4, we examine the sensitivity of the ratio estimator to deviations from the underlying assumptions.
-We first consider the case in which the spatial distribution of the survey respondents is not representative, and we examine a more general estimator in which sampling weights adjust for differential nonresponse.
-We then consider the case in which the Poisson process assumption is violated, and we examine a more general model in which the arrival of street vendors follows a pure birth process, of which the Poisson and Yule processes are special cases.
-The Yule process may be more appropriate if street vendors cluster within markets.
-In both cases, we rederive the standard errors.
+We first consider the scenario in which the spatial distribution of the survey respondents is not representative, and we examine a more general estimator in which sampling weights adjust for differential nonresponse.
+We then consider the scenario in which the Poisson process assumption is violated, and the negative binomial process provides a better approximation.
+For both scenarios, we rederive the standard errors.
 In Section 5, we estimate the number of street vendors in New York City.
 We find approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,400 general merchandise vendors.
-We also find that this estimate is somewhat insensitive to deviations from our underlying assumptions.
-For example, we find that if the spatial distribution of survey respondents is not representative, the estimated number of street vendors in New York City could be as low as 21,000 and as high as 32,000.
-In Section 6, we discuss some implications of our estimates.
-In particular, we find that although the spatial distribution of respondents matches the results of the American Community Survey, the two disagree on the size of New York City’s street vending population.
+We also find that this estimate is somewhat sensitive to deviations from our underlying assumptions.
+For example, we find that if the spatial distribution of survey respondents is not representative, the estimated number of street vendors in New York City could be as low as 17,200 and as high as 41,800.
+In Section 6, we conclude with a discussion.
+We first consider how the theoretical arguments provided in Sections 3 and 4 could be extended to more complex models.
+We then consider some implications of the estimates provided in Section 5.
+We find that although the spatial distribution of respondents is consistent with the results of the American Community Survey, the two disagree on the size of New York City’s street vending population.
 Our estimates suggest the American Community Survey misses the majority of New York City street vendors.
-This finding is important because the American Community Survey is commonly used for policy and advocacy, and our estimates shed light on the broader challenges of measuring the “gig economy.”
-2. Background and Survey Design
+We believe these findings shed light on the broader challenges of measuring the “gig economy.”
+2. Background and Survey Operations
 A street vendor is any individual who sells goods from a mobile vending unit instead of a store.
-We distinguish between the mobile vending unit from which the goods are sold (i.e., the vending establishment) and the individuals who own and/or operate that unit (i.e., the vendors).
+We distinguish between the vending establishment—the mobile vending unit from which the goods are sold—and the vendors—the individuals who own and/or operate that unit.
 Note that an establishment refers to a single vending unit.
@@
 For example, suppose Sally owns the business Sally’s Salads, which consists of two food trucks, each staffed by three employees.
-Then Sally’s Salads counts as one business, two establishments, and seven vendors.
+Then Sally’s Salads counts as one business, two establishments, and seven vendors including Sally.
 We limit our analysis to two types of street vendors: mobile food vendors—vendors who sell food items such as sandwiches, drinks, and cut fruit—and general merchandise vendors—vendors who sell merchandise items such as electronics, clothing, and accessories.
@@
 Most relevant is the fact that a merchandise vendor requires a license to operate legally, and a food vendor requires both a license and a permit.
-The number of food permits and merchandise licenses are limited by law: 5,100 permits and 853 general merchandise licenses are available.
-Mobile food vending licenses are not limited.
-Note that a mobile food vending permit is issued to an individual or business to allow for the sale of food from a specific mobile food vending unit, such as a cart or truck.
-A mobile food vending license authorizes an individual to prepare or serve food from a permitted mobile food vending unit.
-There are several versions of a mobile food vendor permit with varying restrictions.
+The number of food permits and merchandise licenses are limited by law: 5,100 permits and 853 general merchandise licenses are available at the time of our writing.
+Food licenses are not limited.
+Note that a food permit is issued to an individual or business to allow for the sale of food from a specific mobile food vending unit, such as a cart or truck.
+A food license authorizes an individual to prepare or serve food from a permitted mobile food vending unit.
+There are several versions of a food permit with varying restrictions.
 Of the 5,100 permits available, 200 are borough permits that limit vendors to one of the five boroughs; 100 are reserved for veterans or vendors with a disability; 1,000 are seasonal and valid only from April to October; 1,000 are green cart permits that limit vendors to selling fruit, vegetables, plain nuts, and water; and 2,800 are unrestricted.
-Multiple individuals with food vendor licenses can legally operate from one permitted food vending unit.
+Multiple individuals with food licenses can legally operate from one permitted food vending unit.
 Merchandise licenses are renewed annually, while food licenses and permits are renewed biennially.
@@
 Our estimates are limited to food and merchandise vendors operating outside of parks.
-2.2 Survey Design
+2.2 Survey Operations
 The Street Vendor Project (SVP), part of the Urban Justice Center, is a non-profit organization that advocates on behalf of New York City street vendors.
@@
 First Amendment vendors were also eligible.
-There were no limits based on the size of the business, the number of sales, or whether licensed and/or permitted.
+There were no restrictions based on residency, the size of the business, the number of sales, or whether licensed and/or permitted.
 All individuals eligible for aid were invited to complete a survey.
@@
 The number of respondents is listed by neighborhood in Table 1.
-(Neighborhoods with few respondents are grouped together.) A map of the number of respondents by ZIP Code Tabulation Area (ZIP Code) is shown in the top left panel of Figure 1.
+Neighborhoods with few respondents are grouped together.
+A map of the number of respondents by ZIP Code Tabulation Area (ZIP Code) is shown in the top left panel of Figure 1.
 Approximately five percent of vendors did not identify their vending location.
 These vendors are included in the New York City total, but they are excluded from the neighborhood estimates in Table 1 and Figure 1.
-(For this reason, the Respondent and Population columns do not sum exactly to the New York City total.) Veteran merchandise vendors are excluded entirely from Table 1 and Figure 1, although these vendors are reflected in our overall estimate of 23,000 vendors.
+For this reason, the Respondent and Population columns do not sum exactly to the New York City total.
+Veteran merchandise vendors are excluded entirely from Table 1 and Figure 1, although these vendors are reflected in our overall estimate of 23,000 vendors.
 3. Methodology
@@
 We use ratio estimation, which leverages the fact that the number of licenses and permits are limited by local law.
-In Section 3.1, we provide a simple explanation of ratio estimation, which is intended to be accessible to a general readership and highlights our main assumption.
-In Section 3.2, we develop a theoretical justification based on the theory of point processes, which, while more technical, allows us to explicitly state the estimands and underlying assumptions and to derive the standard errors.
-Inference is conducted using increasing domain asymptotics, and the details are outlined in Sections 7.1, 7.2, and 7.3.
+In Section 3.1, we provide a simple explanation of ratio estimation.
+The explanation is intended to be accessible to a general readership and highlights our main assumption.
+In Section 3.2, we develop a theoretical justification based on the theory of point processes, which while more technical, allows us to explicitly state the estimands and underlying assumptions and to derive the standard errors.
+The details are outlined in Sections 7.1, 7.2, and 7.3.
 3.1 Ratio Estimation
@@
 The purpose is to highlight the main assumption.
-A more rigorous derivation is presented in Section 3.2.
 Consider a fixed region $A$, and let $N_{i}(A)$ denote the number of individuals in region $A$ of status $i$.
@@
 We define the ratio estimator as
-$\hat{\Lambda}_{0}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}$
+$\hat{\Lambda}_{0}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}.$
 It is called the ratio estimator because it depends on the ratio of random variables $n_{0}(A)\,/\,n_{1}(A)$.
-(The total number of permitted vendors, $N_{1}(A)$, is considered fixed.) The estimator is approximately equal to $N_{0}(A)$ when the response rate, $p$, is approximately the same for each status, i.e.,
-$p\approx\frac{n_{0}(A)}{N_{0}(A)}\approx\frac{n_{1}(A)}{N_{1}(A)}$
-in which case $\hat{\Lambda}_{0}(A)$ is the solution for $N_{0}(A)$ in terms of $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(A)$, obtained by cross-multiplication.
+The total number of permitted vendors, $N_{1}(A)$, is considered fixed.
+We formally define the estimand that $\hat{\Lambda}_{0}(A)$ estimates in Section 3.2.
+For now, we note that the estimator is approximately equal to $N_{0}(A)$ when the response rate, $p$, is approximately the same for each status, i.e.,
+$p\approx\frac{n_{0}(A)}{N_{0}(A)}\approx\frac{n_{1}(A)}{N_{1}(A)}.$
+The formula for $\hat{\Lambda}_{0}(A)$ is obtained by using cross-multiplication to solve this equation for $N_{0}(A)$ in terms of $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(A)$.
 It follows that $\hat{\tau}(A)=\hat{\Lambda}_{0}(A)+N_{1}(A)$ is approximately equal to the total $N_{0}(A)+N_{1}(A)$.
@@
 If we further assume that the response rate does not vary by subregion, we can also determine the total number of mobile food vendors by subregion.
-For example, suppose the response rate in subregion $B\subset A$ is also approximately equal to $p$, i.e.,
-$p\approx\frac{n_{1}(A)}{N_{1}(A)}\approx\frac{n_{0}(B)}{N_{0}(B)}$
+For example, suppose the response rate in subregion $B\subseteq A$ is also approximately equal to $p$, i.e.,
+$p\approx\frac{n_{1}(A)}{N_{1}(A)}\approx\frac{n_{0}(B)}{N_{0}(B)}.$
 Then the subregion estimator
-$\tilde{\Lambda}_{0}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}$
+$\tilde{\Lambda}_{0}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)},$
 is the solution for $N_{0}(B)$ in terms of $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(B)$.
@@
 This is the case with the survey data described in Section 2.2.
-We observe $N_{1}(A)$, the number of permits city-wide, but not $N_{1}(B)$, the number of permits within subregion $B$.
-(Recall that the vending locations of the respondents are recorded and thus $n_{0}(B)$ and $n_{1}(B)$ are observed for any $B\subset A$.)
+We observe $N_{1}(A)$, the number of permits citywide, but not $N_{1}(B)$, the number of permits within subregion $B$.
+Recall that the vending locations of the respondents are recorded and thus $n_{0}(B)$ and $n_{1}(B)$ are observed for any $B\subseteq A$.
 The subregion estimator $\tilde{\Lambda}_{0}(B)$ works by replacing the term $N_{1}(B)$ in the ratio estimator $\hat{\Lambda}_{0}(B)$ with
-$\dfrac{N_{1}(A)\,n_{1}(B)}{n_{1}(A)}$
+$\dfrac{N_{1}(A)\,n_{1}(B)}{n_{1}(A)},$
 which is justified when
-$p\approx\frac{n_{1}(A)}{N_{1}(A)}\approx\frac{n_{1}(B)}{N_{1}(B)}$
-This suggests the estimator
+$p\approx\frac{n_{1}(A)}{N_{1}(A)}\approx\frac{n_{1}(B)}{N_{1}(B)}.$
+This suggests the subtotal estimator
 $\tilde{\tau}(B)=\tilde{\Lambda}_{0}(B)+\dfrac{N_{1}(A)\,n_{1}(B)}{n_{1}(A)}=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(A)}$
@@
 Any sufficiently large area can be partitioned into a large number of theoretically vendable locations.
-Whether a vendor is located within a partition has a vanishingly small probability, such that the number of vendors in that area is well approximated by a Poisson distribution.
+Whether a vendor is located within a partition has a vanishingly small probability such that the number of vendors in that area is well approximated by a Poisson distribution.
 This approximation is accurate even under weak dependence (Freedman 1974).
 Other justifications are possible.
-For example, see Section 4.2 for a justification of the Poisson process based on the theory of birth processes.
+See Section 4.2 for a justification of the Poisson process based on the theory of birth processes.
+Let $\lambda_{i}(x)$ denote the intensity of the process $\Pi_{i}$ at location $x$ so that the number of individuals in any subregion $B\subseteq A$, $N_{i}(B)$, is distributed Poisson with mean measure $\Lambda_{i}(B)=\int_{B}\,\lambda_{i}(x)\,dx$, i.e.,
+$N_{i}(B)\sim\text{Poisson}\left(\int_{B}\,\lambda_{i}(x)\,dx\right).$
+We assume each individual responds independently to the survey with probability $p_{i}(x)>0$ such that by Campbell’s Theorem (Kingman 1992), the respondents constitute a “thinned” inhomogeneous Poisson process, $\pi_{i}$, with intensity $\lambda_{i}(x)\,p_{i}(x)$ at location $x$.
+The number of respondents, $n_{i}(B)$, is then distributed Poisson with mean $\int_{B}\,\lambda_{i}(x)\,p_{i}(x)\,dx$, i.e.,
+$n_{i}(B)\sim\text{Poisson}\left(\int_{B}\,\lambda_{i}(x)\,p_{i}(x)\,dx\right).$
+The data are the locations of the respondents by status (i.e., the realizations of $\pi_{i}$) and the number of vendors in region $A$ with a permit, $N_{1}(A)$.
+In other words, we assume $n_{0}(B)$ and $n_{1}(B)$ are observed for any $B\subseteq A$, while $N_{0}(B)$ and $N_{1}(B)$ are unobserved with the exception of $N_{1}(A)$.
+Our goal is to estimate $\mathbb{E}\left[N_{0}(B)\,|\,N_{1}(A)\right]$, the expected number of street vendors in subregion $B\subseteq A$ without a permit, as well as $\tau(B)=\mathbb{E}\left[N_{0}(B)+N_{1}(B)\,|\,N_{1}(A)\right]$, the expected total with or without a permit.
+We always condition on $N_{1}(A)$ to reflect the fact that while the number of vendors of either status in any subregion $B\subsetneq A$ is random, the total number of vendors with permits citywide is fixed by law.
+However, we sometimes suppress the fact that $N_{1}(A)$ is fixed in our notation.
+For example, $N_{0}(B)$ is assumed to be independent of $N_{1}(A)$ so we often write $\mathbb{E}\left[N_{0}(B)\,|\,N_{1}(A)\right]=\Lambda_{0}(B)$ and $\tau(B)=\Lambda_{0}(B)+\mathbb{E}\left[N_{1}(B)\,|\,N_{1}(A)\right]$.
+In Section 3.1, we described the main assumption of ratio estimation: the status and spatial distribution of the respondents is representative.
+We now make this statement precise.
+We assume that $p_{i}(x)$ can be decomposed into a constant that does not depend on $i$ or $x$ plus a spatially varying error term, $p_{i}(x)=p\,+\,\epsilon_{i}(x)$ and that the error term is orthogonal to the corresponding intensity.
+i.e.,
+$\langle\lambda_{i},\epsilon_{i}\rangle=\int_{B}\lambda_{i}(x)\,\epsilon_{i}(x)\,dx=0.$
+It follows that
+$n_{i}(B)\sim\text{Poisson}\left(p\int_{B}\lambda_{i}(x)\,dx\right)$
+so that by conditioning on $N_{1}(A)$, we arrive at the following probability model for the number of respondents with and without permits
+$n_{0}(B)\sim$ $\ \text{Poisson}\big(p\,\Lambda_{0}(B)\big)$
+$n_{1}(B)\,|\,N_{1}(A)\sim$ $\ \text{Binomial}\big(N_{1}(A),\,p\,q(B)\big)$
+where
+$q(B)=\frac{\Lambda_{1}(B)}{\Lambda_{1}(A)}.$
+The maximum likelihood estimates for $\Lambda_{0}(A)$ and $p$ are
+$\hat{\Lambda}_{0}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\qquad\text{and}\qquad\hat{p}=\dfrac{n_{1}(A)}{N_{1}(A)}.$
+The estimator $\hat{\Lambda}_{0}(A)$ is asymptotically normal with mean $\Lambda_{0}(A)$ and standard error
+$\hat{\text{SE}}[\hat{\Lambda}_{0}(A)]=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(A)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}.$
+The maximum likelihood estimate for the total $\tau(A)$ is $\hat{\tau}(A)=\hat{\Lambda}_{0}(A)+N_{1}(A)$, which is also asymptotically normal with mean $\tau(A)$ and the same standard error as $\hat{\Lambda}_{0}(A)$.
+See Section 7.2 for details.
+The maximum likelihood estimate for $\Lambda_{0}(B)$ is
+$\tilde{\Lambda}_{0}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)},$
+which is asymptotically normal with mean $\Lambda_{0}(B)$ and standard error
+$\tilde{\text{SE}}[\tilde{\Lambda}_{0}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}.$
+Finally, the maximum likelihood estimate for the subregion total $\tau(B)$ is
+$\tilde{\tau}(B)=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(A)},$
+which is also asymptotically normal with mean $\tau(B)$ and standard error
 […]
 […]
-[…]
-[…]
-The data are the locations of the respondents by status (i.e., the realizations of $\pi_{i}$) and the number of vendors in region $A$ with a permit, $N_{1}(A)$.
-In other words, we assume $n_{0}(B)$ and $n_{1}(B)$ are observed for any $B\subseteq A$, while $N_{0}(B)$ and $N_{1}(B)$ are unobserved with the exception of $N_{1}(A)$.
-Our goal is to estimate the expected number of street vendors in subregion $B\subseteq A$ without a permit,
-[…]
-and the expected total with or without a permit,
-$\tau(B)=\Lambda_{0}(B)+\mathbb{E}\left[N_{1}(B)\,|\,N_{1}(A)\right]$
-We condition on $N_{1}(A)$ to reflect the fact that while the number of vendors of either status in any subregion $B\subsetneq A$ is random, the total number of vendors with permits citywide is fixed by law.
-In Section 3.1, we described the main assumption of ratio estimation: that the status and spatial distribution of the respondents is representative.
-We now make this statement precise.
-We assume that $p_{i}(x)$ can be decomposed into a constant that does not depend on $i$ or $x$ plus a spatially varying error term, $p_{i}(x)=p\,+\,\epsilon_{i}(x)$, and the error term is orthogonal to the corresponding intensity.
-i.e.,
-[…]
-It follows that
-[…]
-so that by conditioning on $N_{1}(A)$, we arrive at the following probability model for the number of respondents with and without permits
-$n_{0}(B)\sim$ $\ \text{Poisson}\big{(}p\,\Lambda_{0}(B)\big{)}$
-$n_{1}(B)\,|\,N_{1}(A)\sim$ $\ \text{Binomial}\big{(}N_{1}(B),\,p\,q(B)\big{)}$
-where
-$q(B)=\frac{\Lambda_{1}(B)}{\Lambda_{1}(A)}$
-The maximum likelihood estimates for $\Lambda_{0}(A)$ and $p$ are
-$\hat{\Lambda}_{0}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\qquad\text{and}\qquad\hat{p}=\dfrac{n_{1}(A)}{N_{1}(A)}$
-Using increasing domain asymptotics, $\hat{\Lambda}_{0}(A)$ is asymptotically normal with mean $\Lambda_{0}(A)$ and standard error
-$\hat{\text{SE}}[\hat{\Lambda}_{0}(A)]=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(A)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
-The maximum likelihood estimate for the total $\tau(A)$ is $\hat{\tau}(A)=\hat{\Lambda}_{0}(A)+N_{1}(A)$, which has the same standard error as $\hat{\Lambda}_{0}(A)$. See Section 7.1.2 for details.
-The maximum likelihood estimate for $\Lambda_{0}(B)$ is
-$\tilde{\Lambda}_{0}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}$
-which is asymptotically normal with mean $\Lambda_{0}(B)$ and standard error
-$\tilde{\text{SE}}[\tilde{\Lambda}_{0}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
-Finally, the maximum likelihood estimate for the subregion total $\tau(B)$ is
-$\tilde{\tau}(B)=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(A)}$
-The asymptotic standard error is
-$\tilde{\text{SE}}[\tilde{\tau}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\frac{1}{n_{1}(A)}-\frac{1}{N_{1}(A)}+\dfrac{n_{1}(B)\,n_{1}(A\backslash B)}{n_{1}(A)\,n_{0}(B)^{2}}}$
-where $A\backslash B$ denotes the complement of subregion $B$. See Section 7.1.3 for details.
+Note that while we derive asymptotic results for the expected number of vendors, a prediction interval for $N_{0}(B)$ or $N_{0}(B)+N_{1}(B)$ given $N_{1}(A)$ can also be constructed by noting that the distribution of $N_{i}(B)$ is approximately Poisson with rate $\hat{\Lambda}_{i}(B)$.
 4. Model Validation
@@
 The main assumption is that the status and spatial distribution of the respondents is representative.
-This assumption is necessary to ensure the ratio estimator is consistent.
-A secondary assumption is that the spatial distribution of vendors is well described by a family of inhomogeneous Poisson processes.
-This assumption is necessary to derive the standard errors.
+This assumption is sufficient to ensure the ratio estimator is consistent.
+A secondary assumption is that the spatial distribution of vendors is well approximated by a family of inhomogeneous Poisson processes.
+This assumption is sufficient to derive the standard errors analytically.
 In this section, we relax both assumptions.
@@
 In Section 4.1, we examine a generalization of the ratio estimator in which sampling weights adjust for differential nonresponse.
-In Section 4.2, we model the arrival of street vendors as a pure birth process, and we derive a more general formula for the standard errors in which the amount of extra-Poisson variation can be determined by the number of vendors that cluster within markets.
+In Section 4.2, we model the spatial distribution of respondents using an inhomogeneous negative binomial process, and we derive a more general formula for the standard errors in which extra-Poisson variation or “overdispersion” can be explained by the number of vendors that cluster within markets.
+The details are outlined in Sections 7.4 and 7.5.
 4.1 Representativeness
@@
 If this assumption is violated, then $\hat{\Lambda}_{0}(B)$ may be inconsistent.
-For example, if respondents are more likely to come from locations in which a higher proportion of vendors have permits, then the ratio $n_{0}(B)\,/\,n_{1}(B)$ would not be representative, and $\hat{\Lambda}_{0}(B)$ would likely underestimate $\Lambda_{0}(B)$, even in large samples.
+For example, if respondents are more likely to come from locations in which a higher proportion of vendors have permits, then the ratio $n_{0}(B)\,/\,n_{1}(B)$ would not be representative and $\hat{\Lambda}_{0}(B)$ would likely underestimate $\Lambda_{0}(B)$, even in large samples.
 We think the representativeness assumption is reasonable because the Street Vendor Project canvassed New York City to distribute a large amount of relief aid as discussed in Section 2.
 Nevertheless, to evaluate how the estimates might change when this assumption is violated, we generalize the estimator in Section 3.2 to accommodate spatially varying sampling weights.
-The generalization is useful for determining the sensitivity of our estimates to differential nonresponse as we demonstrate in Section 5.1.
+The generalization allows us to determine the sensitivity of our estimates to differential nonresponse as we demonstrate in Section 5.1.
 Let $w_{0}(x)$ and $w_{1}(x)$ denote two spatially varying weight functions. We define the weighted counts
@@
 where $\pi_{i}\cap B$ refers to the points of the process $\pi_{i}$ that fall in subregion $B\subseteq A$. We also define the weighted estimators for $\Lambda_{0}(A)$ and $p$,
-$\hat{\Lambda}_{0}^{w}(A)=\dfrac{N_{1}(A)\,n^{w}_{0}(A)}{n^{w}_{1}(A)}\qquad\text{and}\qquad\hat{p}^{w}=\frac{n^{w}_{1}(A)}{N_{1}(A)}$
-Under Campbell’s Theorem, the means $\mathbb{E}[n^{w}_{0}(A)]$ and $\mathbb{E}[n^{w}_{1}(A)|N_{1}(A)]$ are
-[…]
-We assume the weight functions $w_{0}$ and $w_{1}$ can be chosen in such a way that the expected value is inversely proportional with the probability of a response. That is,
-$p=\mathbb{E}[w_{0}(x)]\,p_{0}(x)=\mathbb{E}[w_{1}(x)]\,p_{1}(x)$
+$\hat{\Lambda}_{0}^{w}(A)=\dfrac{N_{1}(A)\,n^{w}_{0}(A)}{n^{w}_{1}(A)}\qquad\text{and}\qquad\hat{p}^{w}=\frac{n^{w}_{1}(A)}{N_{1}(A)}.$
+Under Campbell’s Theorem, the expected values $\mathbb{E}[n^{w}_{0}(A)]$ and $\mathbb{E}[n^{w}_{1}(A)\,|\,N_{1}(A)]$ are
+$\mathbb{E}\left[n^{w}_{0}(A)\right]=\int_{A}\mathbb{E}\bigl[w_{0}(x)\bigr]\,p_{0}(x)\,\lambda_{0}(x)\,dx\qquad\text{and}\qquad\mathbb{E}\left[n^{w}_{1}(A)\,|\,N_{1}(A)\right]=\frac{N_{1}(A)}{\Lambda_{1}(A)}\,\int_{A}\mathbb{E}\bigl[w_{1}(x)\bigr]\,p_{1}(x)\,\lambda_{1}(x)\,dx.$
+We assume the weight functions $w_{0}$ and $w_{1}$ can be chosen in such a way that the expected values are inversely proportional to the probability of a response. That is,
+$p=\mathbb{E}[w_{0}(x)]\,p_{0}(x)=\mathbb{E}[w_{1}(x)]\,p_{1}(x).$
+This choice is similar in spirit to Horvitz and Thompson (1952) in that we seek to correct for the influence of $p_{0}(x)$ and $p_{1}(x)$.
+Note that we only require the average weight to be inversely proportional to the probability of a response.
 It follows that $\mathbb{E}[n^{w}_{0}(A)]$ and $\mathbb{E}[n^{w}_{1}(A)\,|\,N_{1}(A)]$ simplify to
-$\mathbb{E}\left[n^{w}_{0}(A)\right]=p\,\Lambda_{0}(A)\qquad\text{and}\qquad\mathbb{E}\left[n^{w}_{1}(A_{i})\,|\,N_{1}(A)\right]=p\,N_{1}(A)$
-and that $\hat{\Lambda}_{0}^{w}(A)$ is approximately normal with mean $\Lambda_{0}(A)$ and standard error
+$\mathbb{E}\left[n^{w}_{0}(A)\right]=p\,\Lambda_{0}(A)\qquad\text{and}\qquad\mathbb{E}\left[n^{w}_{1}(A)\,|\,N_{1}(A)\right]=p\,N_{1}(A)$
+and that $\hat{\Lambda}_{0}^{w}(A)$ is asymptotically normal with mean $\Lambda_{0}(A)$ and standard error
 $\hat{\text{SE}}[\hat{\Lambda}_{0}^{w}(A)]=\dfrac{N_{1}(A)\,n_{0}^{w}(A)}{n_{1}^{w}(A)}\sqrt{\dfrac{\hat{\mu}^{2}_{0}(A)+\hat{\sigma}_{00}(A,A)}{n_{0}^{w}(A)^{2}}+\dfrac{\hat{\mu}_{1}^{2}(A)}{n^{w}_{1}(A)^{2}}+\dfrac{(N_{1}(A)-1)\,\hat{\sigma}_{11}(A,A)}{N_{1}(A)\,n^{w}_{1}(A)^{2}}-\dfrac{1}{N_{1}(A)}-\dfrac{2\,\hat{\sigma}_{01}(A,A)}{n_{0}^{w}(A)\,n_{1}^{w}(A)}}$
 where
-$\hat{\mu}_{k}^{2}(A)=\sum_{x\,\in\,\pi_{k}\cap A}\mathbb{E}\bigl{[}w_{k}^{2}(x)\bigr{]}\qquad\text{and}\qquad\hat{\sigma}_{kl}(A,A)=\sum_{\begin{subarray}{c}x\,\in\,\pi_{k}\cap A\\[1.39998pt] y\,\in\,\pi_{l}\cap A\\[1.39998pt] x\neq y\end{subarray}}\mathbb{C}\bigl{(}w_{k}(x),\,w_{l}(y)\bigr{)}$
-See Section 7.1.4 for details including derivations for the subregion and subtotal estimators.
+$\hat{\mu}_{k}^{2}(A)=\sum_{x\,\in\,\pi_{k}\cap A}\mathbb{E}\bigl[w_{k}^{2}(x)\bigr]\qquad\text{and}\qquad\hat{\sigma}_{kl}(A,A)=\sum_{\begin{subarray}{c}x\,\in\,\pi_{k}\cap A\\[1.63885pt] y\,\in\,\pi_{l}\cap A\\[1.63885pt] x\neq y\end{subarray}}\mathbb{C}\bigl(w_{k}(x),\,w_{l}(y)\bigr).$
+See Section 7.4 for details including derivations for the subregion and subtotal estimators.
 The standard error derived in this section for $\hat{\Lambda}_{0}(A)$ is identical to the standard error $\hat{\Lambda}_{0}(A)$ from Section 3.2 if $w_{0}(x)=w_{1}(x)=1$.
@@
 Note that $\hat{\Lambda}_{0}(A)$ and $\hat{\Lambda}_{0}^{w}(A)$ are related by the relationship
-$\hat{\Lambda}_{0}^{w}(A)=\frac{N_{1}(A)\ n^{w}_{0}(A)}{n^{w}_{1}(A)}=\hat{\Lambda}_{0}(A)\,\frac{n^{w}_{0}(A)\,/\,n_{0}(A)}{n^{w}_{1}(A)\,/\,n_{1}(A)}$
+$\hat{\Lambda}_{0}^{w}(A)=\frac{N_{1}(A)\ n^{w}_{0}(A)}{n^{w}_{1}(A)}=\hat{\Lambda}_{0}(A)\,\frac{n^{w}_{0}(A)\,/\,n_{0}(A)}{n^{w}_{1}(A)\,/\,n_{1}(A)}.$
 It follows that the relative bias depends only on the ratio of the average weights $n^{w}_{0}(A)\,/\,n_{0}(A)$ and $n^{w}_{1}(A)\,/\,n_{1}(A)$.
-For example, if the average permitted vendor is twice as likely to respond as the average unpermitted vendors, $n^{w}_{0}(A)\,/\,n_{0}(A)$ is twice the size of $n^{w}_{1}(A)\,/\,n_{1}(A)$, and $\hat{\Lambda}_{0}(A)$ is twice the size of $\hat{\Lambda}_{0}^{w}(A)$.
+For example, if the average permitted vendor is twice as likely to respond as the average unpermitted vendors, $n^{w}_{0}(A)\,/\,n_{0}(A)$ is twice the size of $n^{w}_{1}(A)\,/\,n_{1}(A)$ and $\hat{\Lambda}_{0}(A)$ is twice the size of $\hat{\Lambda}_{0}^{w}(A)$.
 The estimators $\hat{\tau}(A)$ and $\hat{\tau}^{w}(A)$ are similarly related by the relationship
-$\hat{\tau}^{w}(A)=\frac{N_{1}(A)\left(n^{w}_{0}(A)+n^{w}_{1}(A)\right)}{n^{w}_{1}(A)}=\hat{\tau}(A)\,\frac{\left(n^{w}_{0}(A)+n^{w}_{1}(A)\right)\,/\,\left(n_{0}(A)+n_{1}(A)\right)}{n^{w}_{1}(A)\,/\,n_{1}(A)}$
-The same is true for subregion and subtotal estimators, see Section 7.1.4.
+$\hat{\tau}^{w}(A)=\frac{N_{1}(A)\left(n^{w}_{0}(A)+n^{w}_{1}(A)\right)}{n^{w}_{1}(A)}=\hat{\tau}(A)\,\frac{\left(n^{w}_{0}(A)+n^{w}_{1}(A)\right)\,/\,\left(n_{0}(A)+n_{1}(A)\right)}{n^{w}_{1}(A)\,/\,n_{1}(A)}.$
+The same is true for the subregion and subtotal estimators, see Section 7.4.
 4.2 Overdispersion
@@
 That said, the standard errors derived in Section 3.2 do not necessarily hold when the Poisson assumption is violated.
-We think the Poisson assumption is reasonable because a wide class of data generating processes are well-approximated by the Poisson process under the law of rare events as discussed in Section 3.2.
-Nevertheless, to evaluate how the standard errors might change when this assumption is violated, we consider a general model that describes how street vendors arrived at their present locations.
-Specifically, we consider two families of pure birth processes: one with a constant arrival rate (or “birth rate”) in which the Poisson assumption holds and one with a linear arrival rate in which the Poisson assumption does not hold.
-Suppose vendors serve customers who congregate at fixed locations, which we call markets.
+We think the Poisson assumption is reasonable because a wide class of data generating processes are well-approximated by a Poisson process under the law of rare events as discussed in Section 3.2.
+Nevertheless, to evaluate how the standard errors might change when this assumption is violated, we consider an alternative justification based on the theory of birth processes.
+As in Section 3.2, we only sketch the justification.
+See Ross (2014, chap.
+6) for an extended discussion of the theory of birth processes.
+This second justification asks how street vendors arrived at their present locations, and we consider two possible explanations.
+We first suppose that vendors serve customers who congregate at fixed locations, which we call markets.
 Let $M(B)$ denote the number of markets in any subregion $B\subseteq A\subset\mathbb{R}^{2}$.
-If every market is in a competitive equilibrium such that the expected profit for a new vendor is the same at each market, then the time between vendor arrivals at each market may be well-approximated by an exponential distribution with a constant arrival rate $\theta_{i}$.
-Further suppose each market starts with one vendor at $t=0$.
-Then at the time the survey was conducted ($t=1$), the number of vendors, $N_{i}(B)$, would follow a Poisson distribution with mean $\Lambda_{i}(B)=\theta_{i}\,M(B)$, and the results of Section 3.2 hold.
-Now suppose instead that larger markets grow faster than smaller markets, an empirical phenomenon known as preferential attachment.
-Then the time between vendor arrivals may be better approximated by an exponential distribution with a linear arrival rate, or Yule process.
-That is, let $N^{t}_{i}(B)$ denote the number of vendors in subregion $B$ at time $t$ so that the time between vendor arrivals at each market follows an exponential distribution with rate $\theta_{i}\,N^{t}_{i}(B)$.
-If we assume further that each market started with one vendor at $t=0$, then the number of vendors at the time the survey was conducted ($t=1$), $N_{i}(B)$, would follow a negative binomial distribution with mean $\Lambda_{i}(B)=\exp(\theta_{i})\,M(B)$ and variance $\Lambda_{i}(B)\bigr{(}\Lambda_{i}(B)-M(B)\bigr{)}\,/\,M(B)$.
+We assume markets grow over time as vendors arrive to serve customers.
+If vendor arrivals are well approximated by an exponential distribution with a constant arrival rate or “birth rate”, then the number of vendors at the time of the survey, $N_{i}(B)$, would follow a Poisson distribution as assumed in Section 3.2.
+Suppose instead that larger markets grow faster than smaller markets, an empirical phenomenon known as preferential attachment.
+Then the time between vendor arrivals may be better approximated by an exponential distribution with a linear arrival rate or Yule process.
+In this case, the number of vendors at the time of the survey, $N_{i}(B)$, would follow a negative binomial distribution with mean $\Lambda_{i}(B)$ and variance $\Lambda_{i}(B)\bigr(\Lambda_{i}(B)-M(B)\bigr)\,/\,M(B)$.
 i.e.,
@@
 where we parameterize the negative binomial distribution by its mean and variance for comparison with Section 3.2.
-In this case, $N_{i}(B)$ exhibits extra-Poisson variation, and the results of Section 3.2 do not hold.
-See Ross (2014, chap.
-6) for details.
-To derive the standard errors, suppose a portion of the markets, $p$, were selected for the survey at random so that
-$n_{i}(B)\sim\ \text{Negative Binomial}\left(p\,\Lambda_{i}(B),\ p\,\Lambda_{i}(B)\,\frac{\Lambda_{i}(B)-M(B)}{M(B)}\right)$
+In this case, $N_{i}(B)$ exhibits extra-Poisson variation or “overdispersion” when many vendors cluster within a market and thus the results of Section 3.2 may not hold.
+Note that we define the negative binomial as the total number of trials required to get a predetermined number of successes when sampling with replacement.
+To derive the standard errors, suppose a portion of the markets, $p$, were selected for the survey so that
+$n_{i}(B)\sim\ \text{Negative Binomial}\left(p\,\Lambda_{i}(B),\ p\,\Lambda_{i}(B)\,\frac{\Lambda_{i}(B)-M(B)}{M(B)}\right).$
 The conditional distribution of $n_{1}(B)$ given $N_{1}(A)$ is then
-$n_{1}(B)\,|\,N_{1}(A)\sim\ \text{Negative Hypergeometric}\bigl{(}N_{1}(A)-1,\ M(A)-1,\ p\,q(B)\,M(A)\bigl{)}$
+$n_{1}(B)\,|\,N_{1}(A)\sim\ \text{Negative Hypergeometric}\bigl(N_{1}(A)-1,\ M(A)-1,\ p\,q(B)\,M(A)\bigl).$
 The conditional mean and variance of $n_{1}(A)$ given $N_{1}(A)$ is
 $\mathbb{E}\left[n_{1}(A)\,|\,N_{1}(A)\right]=p\,N_{1}(A)\qquad\text{and}\qquad\mathbb{V}\bigr(n_{1}(A)\,|\,N_{1}(A)\bigr)=u_{1}(A)\,p\,(1-p)\,N_{1}(A)$
-so that under increasing domain asymptotics, the ratio estimator
+so that the ratio estimator
 $\hat{\Lambda}_{0}(A)=\frac{N_{1}(A)\ n_{0}(A)}{n_{1}(A)}$
@@
 where
-$\hat{u}_{0}(B)=\cfrac{\dfrac{N_{1}(A)\ n_{0}(B)}{n_{1}(A)}-M(B)}{M(B)}\qquad\text{and}\qquad u_{1}(A)=\dfrac{N_{1}(A)-M(A)}{M(A)+1}$
-See Section 7.1.5 for details including derivations for the subregion and subtotal estimators. Note that our parameterization makes use of the relationship $q(B)=M(B)\,/\,M(A)$.
+$\hat{u}_{0}(B)=\cfrac{\dfrac{N_{1}(A)\ n_{0}(B)}{n_{1}(A)}-M(B)}{M(B)}\qquad\text{and}\qquad u_{1}(A)=\dfrac{N_{1}(A)-M(A)}{M(A)+1}.$
+See Section 7.5 for details including derivations for the subregion and subtotal estimators. Note that our parameterization makes use of the relationship q(B)=M(B)/M(A)q(B)=M(B)\,/\,M(A).
 The standard error derived in this section for $\hat{\Lambda}_{0}(A)$ is identical to the standard error for $\hat{\Lambda}_{0}(A)$ from Section 3.2 except for the weights $u_{0}$ and $u_{1}$.
-These weights are the result of the extra-Poisson variation induced by preferential attachment—that is, the linear “birth” rate in which the arrival of new vendors is proportional to the number of vendors in the market.
-For example, if there are five vendors in each market on average, then $u_{0}$ and $u_{1}$ are approximately 4 and the standard error of the ratio estimator derived in this section is twice the size of the standard error derived in Section 3.2.
-The same is true for subregion and subtotal estimators, see Section 7.1.5.
+These weights can be explained by preferential attachment—the linear “birth” rate in which the number of new vendors is proportional to the number of vendors already in the market.
+If there are five vendors in each market on average, then $u_{0}$ and $u_{1}$ are approximately 4 and the standard error of the ratio estimator derived in this section is twice the size of the standard error derived in Section 3.2.
+The same is true for subregion and subtotal estimators, see Section 7.5.
 5. Data Analysis
-We apply the results presented in Section 3 and 4 to the data described in Section 2.
-In Section 5.1, we use the ratio estimator described in Section 3.1 and 3.2 to estimate the number of street vendors in New York City.
-In Section 5.2, we use the extensions described in Section 4.1 and 4.2 to deteremine the sensitivity of the estimate to the underlying assumptions.
+We apply the methodology developed in Section 3 and 4 to the data described in Section 2.
+In Section 5.1, we use the ratio estimator from Section 3.1 and 3.2 to estimate the number of street vendors in New York City.
+In Section 5.2, we use the extensions from Section 4.1 and 4.2 to determine the sensitivity of the estimate to the underlying assumptions.
 Sensitivity is assessed using additional data from the NYC Office of Administrative Trials and Hearings (OATH) (accessed 2024-04-03) and the American Community Survey from the U.S.
 Census Bureau (accessed 2022-12-22).
+We also compare our estimates to those obtained using a hierarchical model.
 5.1 Estimation
-According to the survey data described in Section 2.2, $n_{1}(A)=349$ mobile food vendor respondents indicated they had a permit, and $n_{0}(A)=1,051$ indicated they did not.
-Since there are $N_{1}(A)=5,100$ permits available, it follows that the number of vendors without permits is 15,3505,100×1,051/ 349formulae-sequence153505100105134915,350$\approx$ 5,100$\times$ 1,051\,/\,34915 , 350 ≈ 5 , 100 × 1 , 051 / 349.
+According to the survey data described in Section 2.2, $n_{1}(A)=349$ food vendor respondents indicated they had a permit and $n_{0}(A)=1,051$ indicated they did not.
+Since there are $N_{1}(A)=5,100$ permits available, it follows that the number of vendors without permits is $15,350\approx 5,100\times 1,051\,/\,349$.
 Combined with the $5,100$ permitted vendors, the total number of mobile food vendors is approximately $20,500$.
-By similar reasoning—i.e., exploiting the fact that the number of general merchandise licenses is limited for non-veterans—the number of non-veteran merchandise vendors is estimated to be approximately $1,400$.
+By similar reasoning—i.e., exploiting the fact that the number of merchandise licenses is limited for non-veterans—the number of non-veteran merchandise vendors is estimated to be approximately $1,400$.
 When combined with the estimated $1,000$ veteran merchandise vendors (see Section 2.1) and $20,500$ food vendors, the total number of street vendors is approximately $23,000$.
@@
 We also report the margin of error (i.e., two standard errors, half of the width of a 95% confidence interval).
-These estimates can be checked by comparing them to a second, independent assessment conducted by the Street Vendor Project in the Bronx on May 13 and 15, 2022, in which the number of street vendors was documented through on-the-ground observations.
+The margin of error is calculated using the asymptotic approximations from Section 3.2.
+The subregions are chosen so that the estimated number of vendors in each region is sufficiently large that the asymptotic approximations are expected to be accurate.
+We check our estimates by comparing them to a second, independent assessment conducted by the Street Vendor Project in the Bronx on May 13 and 15, 2022, in which the number of street vendors was documented through on-the-ground observations.
 The independent assessment identified 188 total street vendors near Fordham Road (Fordham Road BID and Street Vendor Project 2024).
@@
 We conclude this estimate is consistent with the 188 street vendors counted independently by Fordham Road BID and Street Vendor Project (2024).
-{ThreePartTable}{TableNotes}
-This table provides estimates for the number of mobile food vendors and non-veteran general merchandise vendors.
-The number of veteran general merchandise vendors is approximately 1,000.
-First amendment vendors and vendors in NYC Parks are not included.
 Table 1: Number of Street Vendors
@@
   Total 1,905 21,857 1,941
-[…]
+This table provides estimates for the number of mobile food vendors and non-veteran general merchandise vendors.
+The number of veteran general merchandise vendors is approximately 1,000.
+First amendment vendors and vendors in NYC Parks are not included.
 Figure 1: Location of Street Vendors
@@
 These maps compare the spatial distribution of respondents (left) to the distribution of vendors suggested by administrative data (New York City Office of Administrative Trials and Hearings (OATH), top right) and federal data (American Community Survey (ACS), bottom right).
+The top two maps show the log number of vendors by the ZIP Code of their vending location, and the bottom two maps show the log number of vendors by the PUMA (Public Use Microdata Area) of their residency.
+The top two maps show a similar pattern in which vendors concentrate in North and West Queens and in the Chelsea and Clinton neighborhoods of Manhattan.
+The bottom two maps show a similar pattern in which residencies are more uniformly distributed in North and West Queens, the South Bronx, and Upper Manhattan.
 5.2 Model Validation
@@
 These comparisons are visualized in Figure 1.
-We find the spatial distribution of vending locations largely agree (top panels) as do the spatial distribution of the residencies (bottom panels).
-This suggests that, if there were a region of street vendors missed by the SVP, it was also missed by law enforcement and the ACS.
-We can also combine the data from OATH and ACS with the results from Section 4.1 to assess the sensitivity of the estimates in Section 5.1 to differential nonresponse, as we now demonstrate.
+We find the spatial distribution of vending locations largely agree (top panels) as do the spatial distribution of the residencies (bottom panels) in that discrepancies are relatively minor and explainable by chance.
+We conclude that if there were a region of street vendors missed by the SVP, it was also missed by law enforcement and the ACS.
+We can also combine the data from OATH and ACS with the results from Section 4.1 to assess the sensitivity of the estimates in Section 5.1 to differential nonresponse as we now demonstrate.
 While the difference between these new estimates and the estimates provided in Section 5.1 are not statistically significant at the 95% level, we believe they provide insight into how the estimates in Section 5.1 might theoretically err.
-First, we consider the scenario in which the probability a vendor responds is inversely proportional to the average number of vendors cited for violating street vending laws in the surrounding ZIP Code each day, as reported in the OATH data for the year 2021.
+First, we consider the scenario in which the probability a vendor responds is inversely proportional to the average number of vendors cited for violating street vending laws in the surrounding ZIP Code each day.
 This might be the case if vendors in high enforcement areas are reluctant to interact with survey takers.
@@
 We find the weighted counts for food vendors are $n^{w}_{0}(A)=25.7$ and $n^{w}_{1}(A)=5.24$.
-This suggests there are $\left(n_{0}^{w}(A)+n_{1}^{w}(A)\right)\,/\,\left(n_{0}(A)+n_{1}(A)\right)\,/\left(n_{1}^{w}(A)\,/\,n_{1}(A)\right)=47\%$ more food vendors than estimated in Section 5.1.
+This suggests there are
+$\dfrac{\left(n_{0}^{w}(A)+n_{1}^{w}(A)\right)\,/\,\left(n_{0}(A)+n_{1}(A)\right)}{n_{1}^{w}(A)\,/\,n_{1}(A)}=47\%$
+more food vendors than estimated in Section 5.1.
 A similar argument for merchandise vendors suggests there are 9% more merchandise vendors than estimated in Section 5.1.
 Put together, the weights suggest the number of street vendors in New York City is 40% higher than estimated in Section 5.2.
-The second scenario we consider is one in which the probability a vendor responds in proportional to the average number of vendors cited for violating street vending laws in the surrounding ZIP Code.
+The second scenario we consider is one in which the probability a vendor responds is proportional to the average number of vendors cited for violating street vending laws in the surrounding ZIP Code.
 This might be the case if vendors in high enforcement areas are easier to identify and thus easier to survey.
-Under this scenario, there are 8% fewer food vendors than estimated in Section 5.1 and 9% fewer merchandise vendors, suggesting the number of street vendors in New York City could be 8% lower than estimated in Section 5.2.
-The final two scenarios are identical to the first two scenarios, except we use the number of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers in the surrounding Public Use Microdata Area as estimated from the 2017-2021 U.S.
-Census Bureau (accessed 2022-12-22).
+In this scenario, there are 8% fewer food vendors than estimated in Section 5.1 and 9% fewer merchandise vendors, suggesting the number of street vendors in New York City could be 8% lower than estimated in Section 5.2.
+The final two scenarios are identical to the first two, except we use the number of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers in the surrounding Public Use Microdata Area as estimated from the 2017-2021 U.S.
+Census Bureau (accessed 2022-12-22) (as indicated by the POWPUMA variable).
 The result is less extreme than using the average number of vendors cited.
-If the probability of response is inversely proportional to this estimate, there are 33% more vendors New York City than estimated in Section 5.2.
-If the probability of response is proportional to this estimate, there are 5% fewer vendors.
+If the probability of response is inversely proportional, there are 33% more vendors in New York City than estimated in Section 5.2.
+If the probability of response is proportional, there are 5% fewer vendors.
 Combining these four scenarios suggests a conservative range of between 21,000 and 32,200 street vendors operate in New York City.
-The second assumption is that the spatial distribution of vendors is well described by a family of inhomogeneous spatial Poisson processes.
-We argue that this assumption is reasonable because a wide class of data generating processes are well-approximated by the Poisson process under the law of rare events as discussed in Section 3.2.
-It is possible, however, that vendors cluster within markets as described in Section 4.2.
-The weights $u_{0}$ and $u_{1}$ in Section 4.2 correspond to the relative number of vendors without permits and with permits per market respectively.
+The lower and upper limit of this range are estimates.
+Using the standard deviations in Section 4.1, a conservative 95% confidence interval is 17,200 to 41,800.
+The second assumption is that the spatial distribution of vendors is well approximated by a family of inhomogeneous Poisson processes.
+We argue this assumption is reasonable because a wide class of data generating processes are well-approximated by a Poisson process under the law of rare events as discussed in Section 3.2.
+The fact that many street vendors operate in a few Zip Codes can be explained by socioeconomic factors unique to those Zip Codes and does not necessarily suggest a more complicated model.
+Nevertheless, it is possible that vendors cluster within markets according to a birth process as described in Section 4.2.
+In that case, the spatial distribution of vendors may be better approximated by a family of inhomogeneous negative binomial processes where the weights $u_{0}$ and $u_{1}$ correspond to the relative number of vendors without permits and with permits per market respectively.
 These weights cannot be estimated from the survey data alone.
-However, we believe that $\Lambda_{0}(A)$ and $N_{1}(A)$ are no more than 10 times larger than $M(A)$.
-It follows that $u_{0}$ and $u_{1}$ are less than 9, and thus the standard errors from Section 3.2 and the margins of error from Table 1 are at most 3 times larger if the Poisson assumption is violated and the Yule process holds.
+However, we believe that if there is clustering as described in Section 4.2, $\Lambda_{0}(A)$ and $N_{1}(A)$ are no more than 10 times larger than $M(A)$.
+It follows that $u_{0}$ and $u_{1}$ are less than 9, and thus the standard errors from Section 3.2 and the margins of error from Table 1 are at most 3 times larger.
 We suspect this bound is conservative, however, and the typical market contains only one or two vendors.
+We conclude this section by comparing our results to a hierarchical approach in which both the expected number of vendors and the probability a vendor responds to the survey varies by subregion.
+The hierarchical approach provides an alternative framework for constructing estimates and standard errors, and comparison with the proposed approach allows us to double check the sensitivity of our estimates and standard errors to our underlying assumptions.
+The hierarchical approach still assumes the independent Poisson processes and their thinned counterparts as described in Section 3.2.
+However, we depart from Section 3.2 by allowing $p(B_{i})$, $\Lambda_{0}(B_{i})$, and $\Lambda_{1}(B_{i})$ to vary according to the model
+$\text{logit}\,p(B_{i})\sim$ $\ \text{Normal}(\mu_{p},\,\sigma_{p}^{2})$
+$\text{log}\,\Lambda_{0}(B_{i})\sim$ $\ \text{Normal}(\mu_{0},\,\sigma_{0}^{2})$
+$\text{log}\,\Lambda_{1}(B_{i})\sim$ $\ \text{Normal}(\mu_{1},\,\sigma_{1}^{2}).$
+Note that the hierarchical approach also departs from Sections 4.1 and 4.2 because $\Lambda_{0}(B_{i})$ and $\Lambda_{1}(B_{i})$ now share a common distribution across subregions so that the hyperparameters $\mu_{p}$ and $\sigma_{p}^{2}$ are identified.
+We use the data described in Section 2 and the statistical software Stan to sample from the posterior distribution of the hierarchical model.
+We then use the posterior draws to obtain estimates and uncertainty intervals.
+The posterior mean of the total number of vendors is 21,000 with a 95% credible interval of 18,900 to 23,900, which is consistent with the proposed approach.
+A similar hierarchical model using the negative binomial process introduced in Section 4.2 produces a similar interval.
+See Section 7.6 for details including the Stan code used.
 6. Discussion
-We discuss the implications of our estimates.
-In particular, we assess the completeness of the New York City Office of Administrative Trials and Hearings (OATH) and American Community Survey (ACS) data.
+We conclude with a discussion.
+In Section 6.1, we consider how the theoretical justification provided in Section 3.2 and extended in Sections 4.1 and 4.2 might be further extended to accommodate more complex models.
+In Section 6.2, we consider some implications of the estimates provided in Section 5.
+6.1 Theoretical Extensions
+In Section 3.2, we develop a theoretical justification for ratio estimation based on the theory of point processes.
+We assume the spatial distribution of survey respondents is representative and well-approximated by a set of inhomogeneous Poisson processes.
+Our approach yields simple, closed-form expressions for the estimates and standard errors on which policy can be based.
+In Section 4, we relax both assumptions to determine the sensitivity of these expressions.
+Our relaxations yield additional closed-form expressions for the estimates and standard errors.
+These expressions are more complicated than those derived in Section 3.2.
+While the estimates and standard errors are not identified from the survey data alone, they provide insight into how exactly a violation would change an estimate or standard error, again facilitating policy.
+The results of Section 5 suggest that deviations from our assumptions do not materially change our findings, except in extreme cases that we do not believe hold in practice such as if the response probability were inversely proportional to enforcement.
+However, it is possible our findings would change if we were to relax both assumptions further.
+For example, in Section 4.1 we relax the assumption that the distribution of respondents is representative by generalizing the Poisson process to accommodate probability weights with a known mean and covariance.
+Future work might consider the more general case in which the mean and covariance of the weights are estimated.
+One might model the response probability with additional data using logistic regression from which the mean and covariance of the weights could be estimated.
+The additional uncertainty would inflate the standard errors relative to Section 4.1.
+See Brick and Montaquila (2009, sec.
+3.2) for additional considerations when modeling the response probability.
+For another example, in Section 4.2 we relax the assumption that the spatial distribution is well-approximated by a set of inhomogeneous Poisson processes.
+We consider the possibility that the spatial distribution of vendors is better approximated by a set of inhomogeneous negative binomial process, which produces overdispersion when many vendors cluster within a market.
+Future work might consider even more general point process models.
+For example, one might model the location of vendors as a Neyman-Scott or Cox process, which also produce overdispersion.
+Overdispersion is not inevitable, however, as it is possible that vendors protect their turf, leading to underdispersion.
+It may even be that some vendors benefit from operating near each other, while others benefit from maintaining their distance.
+See Cressie and Wikle (2011, chap.
+4.3) for additional considerations when modeling Cox or dependent processes.
+Finally, we assume that spatial regions are sufficiently large or dense that inference can be achieved using asymptotic arguments.
+Information is borrowed across subregions through estimation of a shared ratio.
+Future work might compare this approach with other methods for sharing information between subregions.
+For example, the literature on small area estimation often uses hierarchical models such as the one specified in Section 7.6 to borrow information across subregions, often called subdomains.
+See Rao (2003) for for additional considerations when estimating counts in small regions.
+One way to augment the hierarchical approach outlined in Section 5.2 is to model the $\mu_{i}$ as a function of auxiliary data, such as those described in that section.
+6.2 Practical Implications
+We consider the completeness of the New York City Office of Administrative Trials and Hearings (OATH) and American Community Survey (ACS) data.
 Our estimates suggest that while both data sets are spatially representative, they miss the majority of street vendors.
@@
 Indeed, a street vendor must be sufficiently visible while on the street in order to sell their items, while the residency of that same vendor may be more easily overlooked.
+Even if the vendor responds to the ACS, they may fail to indicate that they are a street vendor or work in New York City.
+Indeed, approximately 25% of individuals classified as Door-To-Door Sales Workers, News And Street Vendors, And Related Workers with place of work in New York City had their occupation imputed due to missing values as indicated by the occupation allocation flag.
+Even more, 40%, had their work location imputed as indicated by the place of work allocation flag.
+These imputation rates are high but not uncommon.
+Approximately a third of individuals classified as working in New York City had an occupation with a higher allocation rate, while 10% had a higher place of work allocation rate.
 Another explanation is that the Street Vendor Project survey and the ACS may be measuring different populations or concepts.
@@
 Street vending often provides supplemental income, and thus the discrepancy may reflect the broader challenge of studying the “gig economy.” In this case, either estimate may be relevant, depending on the intended use case.
-We believe this application highlights the value of modeling respondents as a point process.
-We believe our approach is relevant for a wide variety of surveys in which units cannot be chosen at random from a sampling frame.
-For example, individuals may be sampled on a beach or at various entrances of a park.
-Future work might explore additional point process models that describe the behavior of the population.
-For example, we have found that when vendors cluster within markets, standard errors become overdispersed.
-However, it is possible that vendors protect their turf, leading to underdispersion.
-It may even be that some vendors benefit from operating near each other, while others benefit from maintaining their distance.